### v1

# Evaluating Perspectival Biases in Cross-Modal Retrieval

Abstract
Multimodal retrieval systems are expected to operate in a semantic space, agnostic to the language or cultural origin of the query. In practice, however, retrieval outcomes systematically reflect perspectival biases: deviations shaped by linguistic prevalence and cultural associations. We study two such biases. First, prevalence bias refers to the tendency to favor entries from prevalent languages over semantically faithful entries in image-to-text retrieval. Second, association bias refers to the tendency to favor images culturally associated with the query over semantically correct ones in text-to-image retrieval. Results show that explicit alignment is a more effective strategy for mitigating prevalence bias. However, association bias remains a distinct and more challenging problem. These findings suggest that achieving truly equitable multimodal systems requires targeted strategies beyond simple data scaling and that bias arising from cultural association may be treated as a more challenging problem than one arising from linguistic prevalence.
Keywords model bias/fairness evaluation $\cdot$ multimodality $\cdot$ multilingual evaluation $\cdot$ language/cultural bias analysis
1 Introduction
As Nietzsche [1] observed, “there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere. Large models inherit this perspectival character through their training data; what they learn to represent depends on the frequency of appearance and co-occurrence structure. As a result, the latent space of such models does not always function as the robust, language-agnostic semantic space we expect. Instead, retrieval outcomes can be skewed, favoring linguistic prevalence or cultural association over true semantic relevance. The effect of such a perspectival character on both image-to-text and text-to-image retrievals is illustrated in Figure 1. Understanding and quantifying these effects is crucial for ensuring consistent retrieval performance across languages and cultures.
Figure 1: Two Forms of Perspectival Biases. (a) Prevalence bias: an image query favors high-resource languages. A retrieval model places English results above semantically equivalent Japanese and Thai captions. (b) Association Bias: A visualized model’s embedding space, demonstrating how a Japanese text query for “necklace” retrieves culturally proximate images (Japanese masks) instead of the semantically correct one (Kenyan necklace).
Multimodal retrieval enables cross-modality search, primarily between text and images. Early models, such as CLIP [2], align vision and language representations through paired supervision. Recent Multimodal Large Language Models (MLLMs) [3, 4] achieve alignment implicitly through large-scale pretraining. Despite these advancements, the critical issue of language and cultural bias in retrieval remains underexplored.
This lack of study is concerning given that state-of-the-art retrievers are trained on web-scale, text-image datasets like LAION [5] and WebLi [6], which are overwhelmingly English-centric. While these datasets are constructed using English alt-text, images with high cultural specificity often retain alt-text in their native languages. As observed in multilingual food datasets [7], items like the Catalan pastry “coca de recapte” are exclusively described in Catalan or Spanish. This might allow models to develop emergent multilingual capabilities, but it also risks introducing a systemic bias where the model learns spurious correlations, preferentially matching images with text from a specific majority or “expected” language.
A key barrier to investigating these biases is the absence of targeted metrics and benchmarks designed to quantify them. To address this gap, we introduce an evaluation framework for both retrieval directions, each capturing a distinct form of perspectival bias. Image-to-text retrieval. In the absence of linguistic cues, retrievals reveal how the prevalence of certain languages in the training data shapes the results. To assess this prevalence bias, we propose the Discounted Language Bias Kullback–Leibler Divergence (DLBKL), inspired by Language Bias Kullback–Leibler Divergence (LBKL) [8], which measures how strongly retrieval relevance depends on language rather than semantics, as shown in Figure 1a. Text-to-image retrieval. When linguistic and cultural cues are present in the query, retrievals reveal the model’s tendency to favor culturally associated visual patterns over semantically aligned ones. We term this association bias and construct a balanced, cross-cultural, and cross-lingual dataset to disentangle semantic relevance from cultural proximity, as shown in Figure 1b.
Using these tools, we conduct an empirical analysis comparing the perspectival biases inherent in retrievers adapted from MLLMs with those trained using explicit cross-lingual alignment techniques [9, 10]. Our findings reveal that models with explicit alignment mechanisms exhibit lower biases, highlighting a critical trade-off between the scale of MLLMs and the fairness of more targeted alignment strategies.
We summarize our contributions as follows: (i) We propose DLBKL, a metric for quantifying prevalence bias in multimodal retrieval within a multilingual candidate pool to assess language fairness. (ii) We introduce a novel vision-language dataset, parallel across culture and language, designed to assess association bias.
2 Related work
2.1 Multimodal Retrievers
Retrieving image information using a text query can be accomplished by two methods: sparse and dense retrieval. Sparse retrieval methods utilize a high-dimensional representation extracted from words in a text query or an image caption [11, 12]. While these methods are fast, they do not understand the semantics of the image as they solely rely on the image caption to represent the content. To handle the challenge of understanding semantic meaning, deep learning-based dense retrieval methods have been developed. For example, CLIP [2] and ALIGN [13] used a dual-encoder architecture specifically trained to connect the semantic meanings of text-image pairs using contrastive loss objectives. Both models were built on similar principles but varied in text-image data size and utilized different text/image encoders. The recent ColQwen [14] and GME [15] model adapts a Large Language Model (LLM) to learn a more intensive semantic connection between text and images, converting them to a multimodal retrieval model.
2.2 Language Bias in Multi-model Retrievers
Language bias in multimodal retrieval refers to performance differences that arise when semantically equivalent queries in different languages yield divergent rankings, often favoring high-resource languages such as English. Prior work frames this along two fairness axes: (i) an individual-level notion, where multilingual variants of the same query should produce similar results, and (ii) a group-level notion, where aggregate retrieval performance should remain balanced across languages [16].
A study on multilingual retrieval benchmark [17] reports uneven performance across languages, with comparatively stronger results on English and other high-resource languages for various modern multimodal retrievers, despite their large scale in data and parameters. For instance, some models exhibit significant variation in NDCG [18] scores across different languages, indicating that retrieval effectiveness is not uniform.
To quantify such disparities, fairness-aware metrics from ranking literature, such as exposure parity [18], have been adapted to language as a protected attribute. More recently, Adewumi et al. [19] surveyed multimodal bias, emphasizing the lack of dedicated language-focused evaluation protocols. Addressing this, Laosaengpha et al. [8] proposed LBKL, a distributional measure of divergence between retrieval results across language variants. Although LBKL was designed for measuring text modality bias, it can technically be extended to multimodal retrieval, enabling a more fine-grained detection of scores across languages. These works highlight several metrics for measuring language bias in multimodal retrievers. However, despite their effectiveness in measuring language bias, none take retrieval rankings into account.
2.3 Multilingual and Cross-Lingual Retrieval Strategies
Two dominant paradigms exist for building multilingual multimodal retrieval systems: holistic end-to-end pre-training and explicit cross-lingual alignment.
Holistic End-to-End Pre-training on Large-Scale Multilingual Data
This approach, anchored by foundation MLLMs like Qwen2.5-VL [3], aims to learn an emergent universal representation from web-scale, mixed-language data. Within this paradigm, some models like GME [15] and the standard ColQwen series [14] are fine-tuned on predominantly English datasets. Others, such as the multilingual ColQwen series [14] and jina-embeddings-v4 [20], intentionally incorporate extensive multilingual data to improve fairness.
Explicit Cross-Lingual Alignment via Knowledge Distillation
This alternative strategy uses knowledge distillation to align text encoders for new languages to a strong, pre-existing English model’s embedding space, such as CLIP’s. This data-efficient method, exemplified by M-CLIP [10], typically requires only parallel text corpora to force non-English embeddings to mimic their English counterparts via a teacher-student setup.
Our work evaluates models representing both paradigms, providing a direct comparison of biases inherent to each approach.
3 Methodology for Evaluating Bias in Multimodal Retrieval
This section outlines the framework developed to investigate perspectival bias in multilingual, multimodal retrieval systems. We first state our guiding research questions and then explain the studies that address these research questions in Sections 3.1 and 3.2. Our investigation centers on two complementary perspectives, corresponding to different retrieval directions, as shown in Figure 2:
  1. RQ1
[…]
  2. RQ2
[…]
Figure 2: Overview of the study. First, in RQ1, we identify language prevalence in image-to-text retrieval by analyzing the language of the retrieved text and comparing it with high-resource languages, as well as medium- and low-resource languages. Second, in RQ2, we identify the association bias using self-preference behavior of the model by retrieving an image with three candidates: semantically relevant, culturally relevant, and a non-relevant candidate.
3.1 RQ1: Image-to-Text Retrieval Study
In this study, we assess the bias arising from linguistic prevalence by examining the discrepancy between an expected “fair” language distribution and the observed one. That is, the discrepancy should be zero if the linguistic prevalence has no effect on the retrieval results and increases as the results deviate from the ideal case. As discussed in Section 2.2, existing work lacks a dedicated metric to quantify such a discrepancy in multimodal retrieval.
As the first step to closing this gap, we apply the Language Bias Kullback–Leibler (LBKL) Divergence proposed by Laosaengpha et al. [8], which measures the divergence between an expected language distribution and the observed distribution in a retrieved list. Given proportions of language A and B in the ground truth ($P_{A}(x),P_{B}(x)$) and in the retrieved set ($Q_{A}(x),Q_{B}(x)$), LBKL is given as:
$\text{LBKL}=\frac{\sum\limits_{i=1}^{q}\left[P_{\text{A}}(x)\log\frac{P_{\text{A}}(x)}{Q_{\text{A}}(x)}+P_{\text{B}}(x)\log\frac{P_{\text{B}}(x)}{Q_{\text{B}}(x)}\right]}{q}$ (1)
While LBKL can be applied to cross-modal retrieval, it is rank-agnostic: deviations at rank 1 are penalized equally to deviations at rank 100. This underestimates the harm in systems that concentrate resource-driven bias near the top ranks.
For the next step, we extend LBKL by introducing the Discounted Language Bias Kullback-Leibler (DLBKL) divergence, which incorporates a logarithmic rank discount inspired by NDCG [18]. We assign a weight $w(i)=1/\log_{2}(i+1)$ to each rank $i$. The rank-weighted proportion for a language $l$ is then:
$Q^{\prime}_{l}(x)=\frac{\sum_{i=1}^{k}w(i)\cdot\mathbb{I}(\text{doc}_{i}\text{ is }l)}{\sum_{i=1}^{k}w(i)}$ (2)
where $\mathbb{I}(\cdot)$ is the indicator function. DLBKL is calculated by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in the LBKL formula. As illustrated in Figure 3, DLBKL penalizes top-ranked disparities more heavily, aligning the metric with user exposure and better capturing the discrepancy between the ideal case and observed one in multimodal retrieval.
Figure 3: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists where high-resource languages dominate the top ranks.
3.2 RQ2: Text-to-Image Retrieval Study
Figure 4: Overview of the XCM dataset creation process, designed to produce a benchmark with parallelism across semantics, cultures, and languages.
To quantify the degree to which models prioritize cultural association over semantic fidelity, a phenomenon we illustrate in Figure 1(b), a benchmark with a parallel structure in its cultural dimension is necessary. To the best of our knowledge, no such benchmark exists, so we make two primary contributions. First, we construct and introduce the Cross-Cultural Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose. Second, we propose the Self-Preference Cultural Bias Score (SP), a new metric for explicitly measuring this form of bias.
3.2.1 The 3XCM Dataset Benchmark
To evaluate association bias, we constructed the 3XCM benchmark [^1] Research release only (CC BY-NC-SA 4.0). Ethical review required for production use.
Available at: https://huggingface.co/datasets/Chula-AI/association_bias_benchmark.
. The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
The image gathering stage, summarized in Figure 4, consisted of three steps:
  • Concept Generation We used Gemini [^2] Version used: gemini-2.5-flash (Released June 17, 2025). to generate a large pool of concepts, which we manually curated to a final set of 138 coarse-grained, culturally-inclusive concepts (e.g., "train", "food"). Each concept is an abstract, semantic category that uses shared properties to group a broad, culturally-inclusive range of entities. The prompt for generating concepts can be found in Appendix A.
  • Concept De-duplication: We use BGE-M3 [21] to de-duplicate concepts based on similarity with a threshold of 0.92.
  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [22] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
  • Image De-duplication: To ensure visual diversity, we performed two-stage de-duplication within each concept. First, near-exact duplicates were removed automatically using an embedding model. Subsequently, three human annotators, following the guidelines in Appendix F, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
Leveraging the collected cultural images, we introduce a novel evaluation paradigm that employs a forced-choice task. This setup is designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association. As illustrated in Figure 5, for a given query (e.g., "food" in Thai), the model is presented with a triplet of image candidates: (i) Semantically Relevant: same concept, different culture (e.g., Nigerian food); (ii) Culturally Relevant: different concept, same culture (e.g., Thai traditional dance); and (iii) Non-Relevant: different concept and culture (e.g., Japanese gas station).
Figure 5: Illustration of association bias evaluation. A Thai text query for “food” is evaluated against three candidates designed to isolate semantic faithfulness vs. cultural relevance.
The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix J and N respectively.
3.2.2 Self-Preference Cultural Bias Score (SP)
With the constructed dataset, we can now measure the discrepancy between the ideal case and the observed one, where bias arising from cultural association may intervene. Ideally, the discrepancy should be zero when image retrieval depends solely on semantic relevance, and it should increase as the model’s preference tends towards images associated with the culture of the query, rather than semantic accuracy. To quantify the discrepancy, we propose a metric called the self-preference cultural bias score (SP), which can be computed as follows:
$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{sem},i},S_{\text{cul},i},S_{\text{non},i})\right)$ (3)
$\text{SP}=\frac{M_{\text{cul}}}{M_{\text{sem}}}$ (4)
where $M_{k}$ is the proportion of times a candidate of type $k$ receives the highest similarity score across $N$ total trials. The candidate type $k$ can be semantically relevant (sem), culturally relevant (cul), or non-relevant (non). The similarity score for candidate type $k$ in trial $i$ is denoted by $S_{k,i}$. The indicator function $\mathbb{I}(\cdot)$ is 1 if the condition is true and 0 otherwise. The SP score (Eq. 4) is then the ratio of cultural wins ($M_{\text{cul}}$) to semantic wins ($M_{\text{sem}}$). In this way, a higher SP score indicates stronger cultural self-preference over semantic faithfulness and thus a greater extent of association bias.
4 Experimental Setup
To answer our research questions, we conducted two main experiments. For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [23]. The dataset offers a balanced multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities. We evaluate models using Accuracy@5, NDCG@10, LBKL@10, and our proposed DLBKL@10. For RQ2, we performed text-to-image retrieval on our newly created XCM benchmark, evaluating models using our proposed SP score.
For both RQ1 and RQ2, we selected a representative suite of models spanning three distinct architectural paradigms, as shown in Table 1:
  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data. We include the original CLIP-L/14 as a powerful baseline, and Chinese-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space. We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
  • MLLM-Based Retrieval Embedders: This modern paradigm adapts large, pre-trained Multimodal Language Models for retrieval. We evaluate several state-of-the-art models, including the ColQwen series (v0.2, 3b-M, 7b-M), GME models (Qwen2-2B, Qwen2-7B), and Jina-E-v4.
Full model identifiers are available in Appendix H.
5 Experimental Results
Our experiments are designed to provide empirical examinations of perspectival biases manifested in image-to-text and text-to-image retrievals.
5.1 Image-to-Text Evaluation (RQ1)
All models exhibit some degree of linguistic prevalence bias. For most MLLM-based models (Jina, ColQwen, GME), the DLBKL score is higher than the LBKL score, as shown in Table 1.This confirms that bias is more pronounced at the top of the ranked list, as these models tend to rank results from medium-to-high resource languages. Results for additional ranks and an example of retrieval result can be found in Appendix D.
Model Acc LBKL DLBKL NDCG
[…]
Vision-Language Contrastive Models
CLIP-L/14 0.509 5.673 5.684 0.290
Chinese-CLIP-L/14 0.355 5.046 5.055 0.207
Cross-lingual Alignment Models
XLM-R-L/14 0.924 0.320 0.333 0.736
XLM-R-B/16plus 0.968 0.110 0.125 0.791
MLLM-Based Retrieval Embedders
ColQwen2.5-3b-M 0.894 0.792 0.817 0.605
ColQwen2.5-7b-M 0.926 0.821 0.849 0.665
ColQwen2.5-v0.2 0.754 3.834 3.867 0.481
GME-Qwen2-2B 0.967 3.121 3.174 0.717
GME-Qwen2-7B 0.979 1.371 1.420 0.770
Jina-E-v4 0.972 0.915 0.951 0.775
Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL. Explicit alignment models (XLM-R) show substantially lower bias.
This phenomenon is visualized in Figure 6, which shows a clear dominance of high-resource languages in the top ranks. This further illustrates the overall disparity in retrieval frequency between language resource tiers as shown in Figure 7.
Figure 6: Distribution of language groups across retrieval ranks. High-resource languages (blue) dominate the top ranks, a bias captured by DLBKL.
Figure 7: Histogram of retrieved language frequencies. MLLM-based models disproportionately retrieve texts from medium-high resource languages (blue) over low-resource ones (orange).
Crucially, the explicit alignment models (XLM-R series) achieve the lowest bias scores by a significant margin, with XLM-R-B/16plus demonstrating near-zero linguistic prevalence bias according to both metrics, while maintaining high retrieval accuracy. This provides strong initial evidence that direct alignment is a more effective strategy for enforcing language fairness than relying on emergent capabilities from large-scale pre-training.
Building on these observations, we note that LBKL and DLBKL quantify distributional bias rather than relevance, and therefore need not correlate with accuracy or NDCG in Table 1. To assess both correctness and fairness, these bias metrics should be interpreted jointly with accuracy (and/or NDCG). Finally, while LBKL/DLBKL capture cross-language imbalance, they do not measure model self-preference (e.g., favoring the query language over others); we operationalize and evaluate that phenomenon with our SP score.
5.2 Text-to-Image Evaluation (RQ2)
Using the proposed XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models. In this evaluation, the semantic win rate ($M_{\text{sem}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias. We observe that the baseline CLIP and Chinese-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
[…]
Vision-Language Contrastive Models
CLIP-L/14 51.24% 40.78% 7.98% 0.80
Chinese-CLIP-L/14 56.39% 31.65% 11.95% 0.56
Cross-lingual Alignment Models
XLM-R-L/14 85.53% 6.84% 7.63% 0.08
XLM-R-B/16plus 87.54% 6.23% 6.24% 0.07
LLM-Based Retrieval Embedders
GME-Qwen2-2B 83.34% 11.64% 5.02% 0.14
GME-Qwen2-7B 84.63% 11.26% 4.11% 0.13
ColQwen2.5-v0.2 82.10% 10.93% 6.97% 0.13
ColQwen2.5-3B-M 83.36% 10.65% 6.00% 0.13
ColQwen2.5-7B-M 84.07% 11.40% 4.53% 0.14
Jina-E-v4 87.56% 7.20% 5.24% 0.08
Table 2: Results on the XCM benchmark for Self-Preference Cultural Bias.
Our culture-specific analysis reveals that this self-preference is a symptom of missing linguistic knowledge, as shown in Figure 8. The CLIP-L/14 model, lacking a robust understanding of non-Latin scripts, defaults to matching cultural origin as a retrieval heuristic. Training on a large Chinese dataset (Chinese-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters. However, this is a shallow fix that fails to generalize to other non-Latin scripts. In contrast, the text-aligned model (XLM-R-L/14) performs well across most languages, with a notable exception for queries in Yoruba (Nigeria). This challenge with low-resource languages persists even in more advanced architectures. For instance, MLLM-based models (Jina-E-v4) employ a LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages. For the vision component, a Vision-Language Model (VLM) is used as the image encoder to improve contextual awareness. However, performance drops for low-resource languages.
Figure 8: Association bias evaluation across four models reveals the limitations of monolingual training. The baseline CLIP (a) shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in Chinese-CLIP (b). In contrast, cross-lingual models like XLM-R (c) and particularly Jina-E-v4 (d) prove far more effective at mitigating this bias and maintaining high semantic relevance across diverse countries.
Figure 9: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) Chinese-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
This behavior is clearly visualized in the UMAP [24] projections of the text embeddings as shown in Figure 9. The baseline CLIP-L/14 model exhibits a fractured embedding space, with non-Latin languages forming distinct clusters far from the main Latin-script cluster. This demonstrates a lack of shared semantic understanding. In the Chinese-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated. In contrast, the explicit alignment model, XLM-R-L/14, successfully unifies the embedding space into a single, language-agnostic cluster, demonstrating a truly shared semantic representation across scripts. The only notable outlier is Yoruba, which was not part of this specific model’s alignment training. The MLLM model, Jina-E-v4, exhibits a similar but distinct pattern: it also forms a single, unified cluster, but the embeddings are more widely dispersed. This suggests a more flexible alignment that may capture finer semantic nuances between languages.
To validate these visual findings numerically, we calculated the silhouette score [25] for each language’s text embeddings. This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix I. This quantitatively reinforces that poor semantic understanding in the text encoder (as visualized by the disparate UMAP clusters) is a key driver of higher cultural association bias.
Both modern MLLM-based models and explicit alignment models drastically reduce the association bias compared to the baselines, achieving SP scores below 0.16. However, neither paradigm consistently outperforms the other on this specific task.
6 Discussion
Our evaluation framework distinguishes between two perspectival biases: prevalence bias, driven by data imbalance, and association bias, arising from learned cultural correlations. Our findings show these are distinct challenges.
Explicit cross-lingual alignment, used by the XLM-R models, is a highly effective strategy, achieving the lowest scores for both prevalence bias (DLBKL) and association bias (SP) by directly enforcing a shared semantic space. While modern MLLMs like Jina-E-v4 also perform well against association bias, the persistence of these issues across all models points to a deeper, unresolved problem: the entanglement of semantic concepts with linguistic and cultural artifacts in the model’s embedding space.
The path forward, therefore, requires a fundamental shift in training strategy. Future work must prioritize training objectives that actively enforce non-association by creating a truly global semantic space. This means designing models to map a semantic query, regardless of its language or cultural origin, to all conceptually relevant images, irrespective of their geographical context. For example, the new method must include a curation process to avoid cross-cultural false negative images being pushed away from their corresponding queries, and utilize data augmentation to help ensure language-agnostic property.
7 Conclusion
In this work, we introduce a framework that distinguishes between two forms of perspectival bias in multimodal retrieval, prevalence bias and association bias, reflecting distinct ways in which a model’s prior shapes its behavior. This conceptual framing is operationalized through our proposed metrics and datasets: DLBKL, which measures rank-aware language prevalence bias, and XCM, which quantifies association bias through cross-cultural image retrieval. Together, these tools enable systematic evaluation of how multimodal large language models inherit and express perspectival biases across languages and cultures.
In image-to-text retrieval, prevalence bias arises when a model favors texts from high-resource languages. This problem is addressed by anchoring other languages to the prevalent ones through cross-lingual alignment. In text-to-image retrieval, association bias arises when a model favors images that are culturally associated with the query language rather than semantically faithful to their content. Such bias cannot be resolved through traditional cross-lingual alignment or by merely exposing the model to a wider range of cultural content during training. Ultimately, our findings call for a more principled approach: one that directly mitigates localized spurious association as a core design principle for models that are not only multilingual but also perform consistently across languages and cultures.
8 Limitations
Our work has several limitations. First, our DLBKL metric measures fairness via distributional parity, not semantic correctness. It therefore cannot distinguish between retrieving irrelevant documents and over-representing a language with relevant ones. Second, the XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures. The benchmark’s coarse-grained semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity.
Acknowledgments
This research has received funding support from the National Science, Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B13F680099] and Chulalongkorn University Artificial Intelligence Institute.
Appendix A Culturally Relevant Concept Identification
To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection. The prompt used in this process is shown in Figure 11 to identify all labels associated with an image. Figure 10.
list 100 concepts that unique and vary in these country including China, India, Japan, saudi arabia, France, German, Brazil,
Kenya, Thailand, USA
like this
{"food":{"China":"Mala Xiang Guo", ..., "Thailand":"Padthai", "USA":"hamburger"}, "costume":{"China":"Hanfu", ...,
"Thailand":"Sabai", "USA":"cowboy"}}
Figure 10: The prompt given to Gemini to generate unique country-specific image concepts.
Appendix B Multi-Concept Detection in Cultural Images
To establish the self-preference cultural bias, the culturally relevant and non-relevant candidate images must not share concepts with the text label. We utilized Gemini2 with the following prompt in Figure 11 to identify all labels associated with an image.
Please classify the following image by assigning them to one or more of the following cultural
categories:
   {category}.
**Comprehensive Output Format (JSON):**
output as JSON for example:
  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
in the categories please order by priority (high to low).
Figure 11: The prompt given to Gemini for multi-label image classification, where the {category} placeholder is dynamically populated with the full list of categories.
Appendix C Language Resources
To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [26] as an approximation. Table 3 and 4 presents the resulting language composition for RQ1 and RQ2, consequently.
Language Type Language ID Distribution (%)
High English en 43.9499
Russian ru 5.7614
German de 5.5691
Medium Japanese ja 4.9152
Chinese-Simpl. zh 4.8778
Spanish es 4.5422
French fr 4.3271
Italian it 2.4060
Portuguese pt 2.3369
Polish pl 1.8744
Dutch nl 1.8083
Indonesian id 1.1759
Turkish tr 1.1274
Czech cs 1.0479
Vietnamese vi 1.0213
Low Korean ko 0.7865
Farsi fa 0.7087
Swedish sv 0.6736
Arabic ar 0.6722
Romanian ro 0.6374
Ukrainian uk 0.6079
Greek el 0.5651
Hungarian hu 0.5082
Danish da 0.4792
Thai th 0.4269
Finnish fi 0.3649
Norwegian no 0.3135
Hebrew he 0.2654
Croatian hr 0.2339
Hindi hi 0.2004
Bengali bn 0.1064
Telugu te 0.0213
Swahili sw 0.0102
Filipino fil 0.0084
Maori mi 0.0014
Cusco Quechua quz 0.0005
Table 3: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ1
Name Full Name Language Resources
USA United States of America English 43.950
UK United Kingdom
AUS Australia
GER Germany German 5.569
CHN China Chinese 4.878
JPN Japan Japanese 4.915
ESP Spain Spanish 4.542
ARG Argentina
FRA France French 4.327
PRT Portugal Portuguese 2.337
BRA Brazil
SAU Saudi Arabia Arabic 0.672
THA Thailand Thai 0.427
IND India Hindi 0.200
KEN Kenya Swahili 0.010
NGA Nigeria Yoruba 0.001
Table 4: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ2
Appendix D Example of Result from Image to Text Retrieval
To further elaborate the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 12. We also provide result of LBKL and DLBKL score at other rank in Table 5.
Figure 12: An Example of Result from Image to Text Retrieval
Model @5 @25 @50 @99
[…]
Vision-Language Contrastive Models
CLIP-L/14 6.904 6.911 4.171 4.182 3.245 3.249 2.658 2.652
Chinese-CLIP-L/14 6.220 6.229 3.793 3.798 3.172 3.168 2.582 2.570
Cross-lingual Alignment Models
XLM-R-L/14 0.939 0.960 0.240 0.246 0.221 0.223 0.214 0.213
XLM-R-B/16plus 0.692 0.713 0.043 0.049 0.030 0.033 0.019 0.019
MLLM-Based Retrieval Embedders
ColQwen2.5-3B-M 2.138 2.164 0.192 0.209 0.114 0.122 0.088 0.089
ColQwen2.5-7B-M 2.373 2.397 0.212 0.232 0.127 0.138 0.083 0.089
ColQwen2.5-v0.2 5.958 5.974 1.375 1.424 0.633 0.676 0.377 0.410
GME-Qwen2-2B 6.014 6.037 0.739 0.804 0.254 0.306 0.140 0.175
GME-Qwen2-7B 3.843 3.875 0.221 0.264 0.094 0.124 0.058 0.077
Jina-E-v4 2.859 2.886 0.157 0.186 0.072 0.093 0.045 0.059
Table 5: Image-to-text retrieval bias on Crossmodal-3600, measured by LBKL and DLBKL at various retrieval depths (k).
Appendix E Language and Rank Frequency Diagram
To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix C, showing both their overall retrieval frequency as shown in Figure 13 and their frequency distribution across ranks as shown in Figure 14.
Figure 13: A language frequency histogram of each language group for all model
Figure 14: A frequency of language group at each rank for all model
Appendix F Annotator Guideline
The guideline we provide to the annotators is to remove duplicates across multiple views. If an image depicts the same scene or object with no meaningful change, keep only one copy. Keep images if there is a significant variation. Allowed differences include time of day (e.g., day vs. night), viewpoint or angle (if the perspective changes enough that visual elements in the image are noticeably different). Minor or trivial variations are not allowed as they would be too similar. This 397 includes slight shifts, crops, or zooms of the same scene.
Appendix G UMAP Analysis for Self-Preference Cultural Bias
To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 15 and 16. The text embeddings cluster strongly by language, a proximity that supersedes semantic content. Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias. While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models. The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster. This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
Figure 15: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CLIP-L/14 model applied to our dataset.
Figure 16: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the Chinese-CLIP-L/14 model applied to our dataset.
Appendix H Full Official Model Name
In this paper, we use aliases for the model names for conciseness; the full names are provided in Table 6.
Alias Used in Paper Full Model Name Parameter
CLIP-L/14 clip-vit-large-patch141 427.6M
Chinese-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
ColQwen2.5-v0.2 ColQwen2.5-v0.23 3814.8M
ColQwen2.5-3B-M ColQwen2.5-3b-multilingual-v1.03 3994.6M
ColQwen2.5-7B-M ColQwen2.5-7b-multilingual-v1.03 8071.1M
GME-Qwen2-2B gme-Qwen2-VL-2B-Instruct4 2209.0M
GME-Qwen2-7B gme-Qwen2-VL-7B-Instruct4 7070.6M
Jina-E-v4 jina-embeddings-v45 3934.7M
XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
The models are based on the following works: 1) Radford et al. [2] for CLIP-L/14; 2) Yang et al. [27] for Chinese-CLIP-L/14; 3) Faysse et al. [14] for ColQwen2 models; 4) Zhang et al. [15] for GME-Qwen2 models; 5) Günther et al. [20] for Jina-E-v4; and 6) Carlsson et al. [10] for XLM-R-VL models.
Table 6: Aliases Used in Paper and Corresponding Full Model Names and Parameters
Appendix I Correlation Analysis for Self-Preference Cultural Bias
We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix G, impacts cross-modal retrieval. To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with self-preference cultural bias score (SP) cultural bias as shown in Table 8. We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores. For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 17. Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder as shown in Table 7.
Figure 17: Comparison of Self-Preference Cultural Bias with Text and Image Silhouette
Model TSC ISC
CLIP-L/14 0.83 0.55
Chinese-CLIP-L/14 -0.26 0.58
XLM-R-VL-B/16 0.98 -0.27
XLM-R-VL-L/14 0.94 0.05
Jina-E-v4 0.86 0.16
GME-Qwen2-2B 0.80 0.09
GME-Qwen2-7B 0.64 0.07
Average 0.68 0.18
Table 7: This table presents a Pearson correlation analysis between model performance and bias. We measure the correlation between association and the quality of data clusters (via Silhouette Score) in both the text embedding space (TSC) and the image embedding space (ISC).
Model Metrics Country
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
CLIP-L/14 SP $\downarrow$ 0.01 0.02 0.02 0.76 2.63 1.94 0.24 0.19 0.34 0.39 0.51 10.71 8.09 15.88 2.04 2.27
TS $\downarrow$ 0.05 0.05 0.05 0.02 0.16 -0.05 -0.01 -0.02 -0.02 -0.01 -0.01 0.25 0.23 0.23 0.03 0.13
IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.04 0.01 0.01 0.00 0.00 0.01 0.05 0.03 0.03 0.02 0.03
Chinese-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
TS $\downarrow$ -0.05 -0.05 -0.05 0.00 0.13 -0.07 -0.03 -0.03 -0.03 -0.01 -0.01 -0.40 0.93 -0.37 0.00 0.03
IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.00 0.00 0.00 0.04 0.05 0.04 0.02 0.01
Jina-E-v4 SP $\downarrow$ 0.02 0.03 0.02 0.05 0.04 0.04 0.04 0.03 0.05 0.04 0.04 0.05 0.02 0.12 0.49 0.93
TS $\downarrow$ -0.01 -0.01 -0.01 -0.02 0.01 0.00 -0.01 -0.02 -0.02 -0.02 -0.02 0.02 0.01 0.00 0.00 0.13
IS $\downarrow$ -0.01 0.00 -0.01 -0.01 0.00 0.01 0.00 0.00 -0.01 -0.01 -0.01 0.02 0.00 0.01 0.01 0.00
XLM-R-L/14 SP $\downarrow$ 0.05 0.04 0.07 0.05 0.04 0.07 0.08 0.08 0.09 0.08 0.07 0.05 0.04 0.04 0.11 0.61
TS $\downarrow$ -0.18 -0.18 -0.18 -0.11 -0.07 -0.10 -0.16 -0.16 -0.16 -0.12 -0.12 -0.07 -0.06 -0.17 -0.14 0.44
IS $\downarrow$ 0.01 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.01 0.01 0.01 0.06 0.04 0.03 0.02 0.03
XLM-R-B/16plus SP $\downarrow$ 0.05 0.04 0.06 0.04 0.04 0.05 0.05 0.06 0.07 0.06 0.06 0.04 0.02 0.05 0.10 0.66
TS $\downarrow$ -0.24 -0.24 -0.24 -0.21 -0.19 -0.19 -0.24 -0.23 -0.23 -0.23 -0.23 -0.19 -0.19 -0.23 -0.23 0.49
IS $\downarrow$ 0.01 0.01 0.00 0.00 0.01 0.01 0.00 0.00 -0.01 0.00 -0.01 0.03 0.02 0.03 0.02 0.00
GME-Qwen2-2B SP $\downarrow$ 0.02 0.03 0.02 0.10 0.04 0.04 0.05 0.05 0.08 0.08 0.09 0.14 0.10 0.24 1.24 2.02
TS $\downarrow$ 0.02 0.02 0.02 0.04 0.02 0.00 0.04 0.01 0.01 0.03 0.03 0.01 0.03 0.00 0.03 0.15
IS $\downarrow$ 0.00 0.00 0.01 0.00 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.02 0.01 0.01
GME-Qwen2-7B SP $\downarrow$ 0.03 0.02 0.01 0.14 0.04 0.05 0.07 0.07 0.07 0.09 0.13 0.14 0.08 0.12 0.62 1.88
TS $\downarrow$ 0.04 0.04 0.04 0.05 0.02 0.02 0.07 0.02 0.02 0.04 0.04 0.09 0.11 0.03 0.04 0.14
IS $\downarrow$ 0.00 0.00 0.01 0.00 0.01 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.01 0.01 0.01
Table 8: Cross-Country and Cross-Model Comparison of Language and Cultural Bias Metrics. This table presents the results for the Self-Preference Cultural Bias score (SP), Text Silhouette (TS), and Image Silhouette (IS) scores across various multimodal retrievers for a selection of countries.
Appendix J Dataset Statistics
The distribution of cultural concepts in the XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
Country Number of Samples
Argentina 771
Australia 721
Brazil 724
China 727
France 760
Germany 744
India 774
Japan 944
Kenya 600
Nigeria 773
Portugal 824
Saudi Arabia 619
Spain 841
Thailand 649
UK 644
USA 609
Average 733
Table 9: Dataset Statistics of XCM dataset per Country
Concepts (A-G)
airlines 24
airport 65
alcohol drink 26
ancient city 112
ancient craft 111
ancient painting 139
animal 100
architecture 107
art 89
artwork 26
bag 20
bakery 47
banknotes 69
bathroom 30
bedroom 77
boat 99
bracelet 62
building 196
bus 92
bus station 56
capital 147
celebrity 66
child 69
Concepts (C-M)
cinema 54
coin 182
combination food 72
congress 127
costume 119
craft 98
dance 145
deep fried food 30
department store 21
dessert 66
devil 43
diningroom 49
doll 131
drink 31
dry heat food 21
embroidery style 133
fashion 57
festival 145
fire station 162
folk tale 38
folklore character 90
food 52
football player 104
Concepts (F-M)
formal uniform 117
fountain style 91
funeral 141
game 76
gas station 63
gathering place 92
ghost 19
graduated uniform 78
hat 95
headwear 106
historical event 89
historical figure 81
historical image 120
hot pot concept 66
hotel 70
house 86
instrument 58
lottery tickets 48
mailbox 81
major mountain range 52
major religious site 97
major river 57
map 32
market 74
Concepts (M-P)
marriage ceremony 95
martial art 96
mask 104
military parade 189
moist heat food 34
museum 99
music band 95
mythical creature 56
mythological figure 68
native inhabitants 163
natural landmark 152
necklace 63
night view 110
older 38
painting 128
pants 30
people 136
poaching food 25
police station 158
popular street food 98
pottery style 150
priest 68
prime minister 86
Concepts (R-S)
religious building 123
restaurant 38
ritual 108
rural dwelling 127
sacred object 101
school 120
series 40
shirt 64
shopping mall 91
singer 56
snack 56
social custom 147
soldier 167
sport 79
stageplay 108
statue 106
street entertainment 111
street sign 81
street vendor cart 32
street view 86
symbolic bird 84
symbolic plant 72
Concepts (T-Z)
tattoo style 59
taxi 102
tea culture 85
textile pattern 56
tourist attraction 130
toy 66
train 116
train station 128
tree 94
tv program 43
unique art form 119
unique cuisine trait 43
unique food ingredient 22
unique natural phenomenon 102
unique transportation 75
university 136
wall painting 65
warrior 79
weapon 75
wedding 108
writing character 15
zoo 79
Table 10: Distribution of Concepts and Image Counts
Model Metrics Country
USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
CLIP-L/14 $M_{\text{sem}}(\%)\uparrow$ 95.73 94.57 94.73 52.42 25.17 31.07 75.53 78.00 69.65 65.78 61.33 7.75 10.48 5.56 27.83 24.19
$M_{\text{cul}}(\%)\downarrow$ 1.31 2.02 2.22 39.65 66.16 60.34 18.29 15.10 23.61 25.85 31.35 83.04 84.75 88.24 56.67 54.85
$M_{\text{non}}(\%)\downarrow$ 2.96 3.42 3.05 7.93 8.67 8.59 6.18 6.90 6.74 8.37 7.32 9.21 4.78 6.20 15.50 20.96
Chinese-CLIP -L/14 $M_{\text{sem}}(\%)\uparrow$ 94.25 89.44 90.57 40.05 93.40 84.20 65.79 61.95 60.83 56.55 57.04 14.38 21.88 10.47 27.00 25.10
$M_{\text{cul}}(\%)\downarrow$ 3.28 5.75 4.44 44.62 2.48 10.92 21.97 24.73 26.07 31.80 29.56 70.11 55.78 73.00 53.83 56.02
$M_{\text{non}}(\%)\downarrow$ 2.46 4.81 4.99 15.32 4.13 4.88 12.24 13.32 13.10 11.65 13.40 15.51 22.34 16.54 19.17 18.89
Jina-E-v4 $M_{\text{sem}}(\%)\uparrow$ 95.89 95.65 95.98 92.47 94.22 92.79 94.34 93.22 91.70 91.99 94.20 91.60 95.38 86.30 56.50 36.74
$M_{\text{cul}}(\%)\downarrow$ 1.97 2.48 1.66 4.97 3.44 3.92 4.08 3.09 4.80 4.13 3.73 4.68 2.00 9.95 27.83 34.15
$M_{\text{non}}(\%)\downarrow$ 2.13 1.86 2.36 2.55 2.34 3.29 1.58 3.69 3.50 3.88 2.07 3.72 2.62 3.75 15.67 29.11
XLM-R-L/14 $M_{\text{sem}}(\%)\uparrow$ 89.66 91.15 89.46 90.32 92.30 89.40 87.63 85.37 86.12 87.86 87.98 90.95 93.07 89.41 80.33 40.49
$M_{\text{cul}}(\%)\downarrow$ 4.11 4.04 5.83 4.44 4.13 6.26 6.58 6.54 7.39 7.16 5.80 4.85 3.54 3.75 8.50 24.71
$M_{\text{non}}(\%)\downarrow$ 6.24 4.81 4.72 5.24 3.58 4.35 5.79 8.09 6.49 4.98 6.22 4.20 3.39 6.85 11.17 34.80
XLM-R-B /16Plus $M_{\text{sem}}(\%)\uparrow$ 91.95 91.30 90.71 93.95 94.50 91.62 90.66 88.47 87.81 90.05 89.36 92.57 95.22 91.09 83.17 40.88
$M_{\text{cul}}(\%)\downarrow$ 4.11 3.88 5.13 4.03 3.58 4.56 4.34 5.47 5.97 5.70 5.52 3.88 2.00 4.91 8.33 26.78
$M_{\text{non}}(\%)\downarrow$ 3.94 4.81 4.16 2.02 1.93 3.82 5.00 6.06 6.23 4.25 5.11 3.55 2.77 4.01 8.50 32.34
GME-Qwen2 -2B-Instruct $M_{\text{sem}}(\%)\uparrow$ 96.06 95.34 96.95 87.77 94.09 93.43 92.76 90.49 88.59 90.05 88.95 84.49 89.06 76.87 37.17 25.87
$M_{\text{cul}}(\%)\downarrow$ 2.30 2.48 1.66 8.47 3.58 3.92 4.74 4.64 7.26 7.04 7.87 11.63 8.63 18.48 46.00 52.26
$M_{\text{non}}(\%)\downarrow$ 1.64 2.17 1.39 3.76 2.34 2.65 2.50 4.88 4.15 2.91 3.18 3.88 2.31 4.65 16.83 21.86
GME-Qwen2 -7B-Instruct $M_{\text{sem}}(\%)\uparrow$ 95.40 95.34 96.53 84.68 95.05 92.47 90.39 90.73 90.27 88.35 85.50 85.46 90.91 85.92 55.17 29.62
$M_{\text{cul}}(\%)\downarrow$ 2.46 2.33 0.97 11.69 3.44 4.77 6.45 5.95 6.10 7.77 11.46 11.63 7.09 10.34 34.00 55.76
$M_{\text{non}}(\%)\downarrow$ 2.13 2.33 2.50 3.63 1.51 2.76 3.16 3.33 3.63 3.88 3.04 2.91 2.00 3.75 10.83 14.62
Table 11: Cross-Country and Cross-Model Comparison of Win Percentages. This table presents the results for the Semantically Relevant, Culturally Relevant, and Non-Relevant Win Percentages across various multimodal retrievers for a selection of countries.
Appendix K Computational Resource
The experiment is performed with a single A100 GPU for approximately 3 gpu hours for each model or 54 hours in total with library version of colpali-engine 0.3.13.dev1+g9bee9b2b7, transformers 4.53.3 for most experiment except, GME models are inferenced under transformers 4.51.3
Appendix L Authoring and Implementation Tools
In preparing this manuscript, we utilized several generative large language models. For language editing and stylistic refinement, we employed Google’s Gemini 2.5-flash, along with models from xAI’s Grok family (e.g., Grok-3 Expert and Fast variants). For assistance with code implementation, scripting, and debugging, we used a model from Anthropic’s Claude series (e.g., Claude 4.0 Sonnet).
Appendix M Detailed Results
The full details of RQ2 experiment including all win rate of all models are illustrated in the Table 11.
Appendix N 3XCM Dataset Benckmark Samples
This research provides an association evaluation benchmark and image metadata. Examples of the benchmark and image metadata are shown in Figure 18 and Figure 19, respectively.
Figure 18: Examples of 3XCM dataset benchmark
Figure 19: Metadata for image of 3XCM dataset benchmark

### v3

@@
 In practice, however, retrieval outcomes systematically reflect perspectival biases: deviations shaped by linguistic prevalence and cultural associations.
-We study two such biases.
-First, prevalence bias refers to the tendency to favor entries from prevalent languages over semantically faithful entries in image-to-text retrieval.
-Second, association bias refers to the tendency to favor images culturally associated with the query over semantically correct ones in text-to-image retrieval.
-Results show that explicit alignment is a more effective strategy for mitigating prevalence bias.
-However, association bias remains a distinct and more challenging problem.
-These findings suggest that achieving truly equitable multimodal systems requires targeted strategies beyond simple data scaling and that bias arising from cultural association may be treated as a more challenging problem than one arising from linguistic prevalence.
-Keywords model bias/fairness evaluation $\cdot$ multimodality $\cdot$ multilingual evaluation $\cdot$ language/cultural bias analysis
+We introduce the Cross-Cultural, Cross-Modal, Cross-lingual Multimodal (3XCM) benchmark to isolate these effects.
+Results from our studies indicate that, for image-to-text retrieval, models tend to favor entries from prevalent languages over those that are semantically faithful.
+For text-to-image retrieval, we observe a consistent “tugging effect” in the joint embedding space between semantic alignment and language-conditioned cultural association.
+When semantic representations are insufficiently resolved, particularly in low-resource languages, similarity is increasingly governed by culturally familiar visual patterns, leading to systematic association bias in retrieval.
+Our findings suggest that achieving equitable multimodal retrieval necessitates targeted strategies that explicitly decouple language from culture, rather than relying solely on broader data exposure.
+This work highlights the need to treat linguistic and cultural biases as distinct, measurable challenges in multimodal representation learning.
 1 Introduction
-As Nietzsche [1] observed, “there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere.
-Large models inherit this perspectival character through their training data; what they learn to represent depends on the frequency of appearance and co-occurrence structure.
-As a result, the latent space of such models does not always function as the robust, language-agnostic semantic space we expect.
-Instead, retrieval outcomes can be skewed, favoring linguistic prevalence or cultural association over true semantic relevance.
-The effect of such a perspectival character on both image-to-text and text-to-image retrievals is illustrated in Figure 1.
-Understanding and quantifying these effects is crucial for ensuring consistent retrieval performance across languages and cultures.
+As nietzsche1887genealogy observed, there is only a perspective seeing, only a perspective knowing”; put differently, there is no view from nowhere.
+Large models inherit this perspectival character from training data, where representations depend on frequency and co-occurrence.
+Consequently, their latent spaces deviate from the expected robust, language-agnostic semantics, skewing retrieval toward linguistic prevalence or cultural associations over true relevance.
+Figure 1 illustrates this in image-to-text and text-to-image retrieval, highlighting the need to quantify such effects for consistent cross-lingual, cross-cultural performance.
 Figure 1: Two Forms of Perspectival Biases.
 (a) Prevalence bias: an image query favors high-resource languages.
-A retrieval model places English results above semantically equivalent Japanese and Thai captions.
-(b) Association Bias: A visualized model’s embedding space, demonstrating how a Japanese text query for “necklace” retrieves culturally proximate images (Japanese masks) instead of the semantically correct one (Kenyan necklace).
-Multimodal retrieval enables cross-modality search, primarily between text and images.
-Early models, such as CLIP [2], align vision and language representations through paired supervision.
-Recent Multimodal Large Language Models (MLLMs) [3, 4] achieve alignment implicitly through large-scale pretraining.
-Despite these advancements, the critical issue of language and cultural bias in retrieval remains underexplored.
-This lack of study is concerning given that state-of-the-art retrievers are trained on web-scale, text-image datasets like LAION [5] and WebLi [6], which are overwhelmingly English-centric.
-While these datasets are constructed using English alt-text, images with high cultural specificity often retain alt-text in their native languages.
-As observed in multilingual food datasets [7], items like the Catalan pastry “coca de recapte” are exclusively described in Catalan or Spanish.
-This might allow models to develop emergent multilingual capabilities, but it also risks introducing a systemic bias where the model learns spurious correlations, preferentially matching images with text from a specific majority or “expected” language.
-A key barrier to investigating these biases is the absence of targeted metrics and benchmarks designed to quantify them.
-To address this gap, we introduce an evaluation framework for both retrieval directions, each capturing a distinct form of perspectival bias.
-Image-to-text retrieval.
-In the absence of linguistic cues, retrievals reveal how the prevalence of certain languages in the training data shapes the results.
-To assess this prevalence bias, we propose the Discounted Language Bias Kullback–Leibler Divergence (DLBKL), inspired by Language Bias Kullback–Leibler Divergence (LBKL) [8], which measures how strongly retrieval relevance depends on language rather than semantics, as shown in Figure 1a.
-Text-to-image retrieval.
-When linguistic and cultural cues are present in the query, retrievals reveal the model’s tendency to favor culturally associated visual patterns over semantically aligned ones.
-We term this association bias and construct a balanced, cross-cultural, and cross-lingual dataset to disentangle semantic relevance from cultural proximity, as shown in Figure 1b.
-Using these tools, we conduct an empirical analysis comparing the perspectival biases inherent in retrievers adapted from MLLMs with those trained using explicit cross-lingual alignment techniques [9, 10].
-Our findings reveal that models with explicit alignment mechanisms exhibit lower biases, highlighting a critical trade-off between the scale of MLLMs and the fairness of more targeted alignment strategies.
-We summarize our contributions as follows: (i) We propose DLBKL, a metric for quantifying prevalence bias in multimodal retrieval within a multilingual candidate pool to assess language fairness.
-(ii) We introduce a novel vision-language dataset, parallel across culture and language, designed to assess association bias.
-2 Related work
-2.1 Multimodal Retrievers
-Retrieving image information using a text query can be accomplished by two methods: sparse and dense retrieval.
-Sparse retrieval methods utilize a high-dimensional representation extracted from words in a text query or an image caption [11, 12].
-While these methods are fast, they do not understand the semantics of the image as they solely rely on the image caption to represent the content.
-To handle the challenge of understanding semantic meaning, deep learning-based dense retrieval methods have been developed.
-For example, CLIP [2] and ALIGN [13] used a dual-encoder architecture specifically trained to connect the semantic meanings of text-image pairs using contrastive loss objectives.
-Both models were built on similar principles but varied in text-image data size and utilized different text/image encoders.
-The recent ColQwen [14] and GME [15] model adapts a Large Language Model (LLM) to learn a more intensive semantic connection between text and images, converting them to a multimodal retrieval model.
-2.2 Language Bias in Multi-model Retrievers
-Language bias in multimodal retrieval refers to performance differences that arise when semantically equivalent queries in different languages yield divergent rankings, often favoring high-resource languages such as English.
-Prior work frames this along two fairness axes: (i) an individual-level notion, where multilingual variants of the same query should produce similar results, and (ii) a group-level notion, where aggregate retrieval performance should remain balanced across languages [16].
-A study on multilingual retrieval benchmark [17] reports uneven performance across languages, with comparatively stronger results on English and other high-resource languages for various modern multimodal retrievers, despite their large scale in data and parameters.
-For instance, some models exhibit significant variation in NDCG [18] scores across different languages, indicating that retrieval effectiveness is not uniform.
-To quantify such disparities, fairness-aware metrics from ranking literature, such as exposure parity [18], have been adapted to language as a protected attribute.
-More recently, Adewumi et al.
-[19] surveyed multimodal bias, emphasizing the lack of dedicated language-focused evaluation protocols.
-Addressing this, Laosaengpha et al.
-[8] proposed LBKL, a distributional measure of divergence between retrieval results across language variants.
-Although LBKL was designed for measuring text modality bias, it can technically be extended to multimodal retrieval, enabling a more fine-grained detection of scores across languages.
-These works highlight several metrics for measuring language bias in multimodal retrievers.
-However, despite their effectiveness in measuring language bias, none take retrieval rankings into account.
-2.3 Multilingual and Cross-Lingual Retrieval Strategies
-Two dominant paradigms exist for building multilingual multimodal retrieval systems: holistic end-to-end pre-training and explicit cross-lingual alignment.
-Holistic End-to-End Pre-training on Large-Scale Multilingual Data
-This approach, anchored by foundation MLLMs like Qwen2.5-VL [3], aims to learn an emergent universal representation from web-scale, mixed-language data.
-Within this paradigm, some models like GME [15] and the standard ColQwen series [14] are fine-tuned on predominantly English datasets.
-Others, such as the multilingual ColQwen series [14] and jina-embeddings-v4 [20], intentionally incorporate extensive multilingual data to improve fairness.
-Explicit Cross-Lingual Alignment via Knowledge Distillation
-This alternative strategy uses knowledge distillation to align text encoders for new languages to a strong, pre-existing English model’s embedding space, such as CLIP’s.
-This data-efficient method, exemplified by M-CLIP [10], typically requires only parallel text corpora to force non-English embeddings to mimic their English counterparts via a teacher-student setup.
-Our work evaluates models representing both paradigms, providing a direct comparison of biases inherent to each approach.
-3 Methodology for Evaluating Bias in Multimodal Retrieval
+A model places English results above semantically equivalent Japanese and Thai captions.
+(b) Association Bias: Given a Japanese text query for “necklace”, a model places a culturally proximate image (Japanese masks) closer to the query than a semantically correct one (Kenyan necklace).
+Multimodal retrieval aligns text and images via paired supervision in early models like CLIP [DBLP:journals/corr/abs-2103-00020] or implicit pretraining in recent Multimodal Large Language Models (MLLMs) [bai2025qwen25vltechnicalreport, comanici2025gemini].
+Yet, language and cultural biases remain underexplored, especially given English-centric datasets like LAION [schuhmann2022laion5bopenlargescaledataset] and WebLi [chen2023pali].
+While culturally specific images may carry native-language alt-text, e.g., Catalan descriptions for “coca de recapte” [olondriz2021foodi].
+This fosters emergent multilingualism but risks spurious correlations, favoring “expected” languages.
+A major barrier is the lack of targeted benchmarks and metrics.
+To address this, we develop an evaluation framework isolating two perspectival biases across retrieval directions.
+For image-to-text (lacking linguistic cues), we assess prevalence bias via the proposed Discounted Language Bias Kullback–Leibler Divergence (DLBKL), extending LBKL [laosaengpha2025mitigatinglanguagebiascrosslingual] to penalize high-resource languages dominating top ranks (Figure 1a).
+For text-to-image (with linguistic/cultural cues), we measure association bias using a parallel, cross-cultural dataset to disentangle semantic fidelity from cultural proximity (Figure 1b).
+Extending this, we test the ability of explicit cultural descriptors (CDs) to override association bias, examining the “tugging effect” where models struggle to prioritize CDs (e.g., “Japanese train” in Thai) over implicit query-language associations.
+Using these tools, we compare biases in MLLM-adapted retrievers against those with explicit cross-lingual alignment [chen-etal-2023-mclip, carlsson-etal-2022-cross].
+Findings expose biases in both directions, including CDs’ limited efficacy: models follow instructions but fall back to language-linked visuals when cultural semantics are unresolved.
+Our contributions: (i) DLBKL as a supplementary, rank-sensitive extension to LBKL for measuring prevalence bias in multilingual pools; (ii) the 3XCM benchmark, parallel across cultures/languages for association bias; (iii) analysis of CDs’ impact on retrieval, revealing persistent biases despite explicit guidance.
+2 Related Works
+2.1 Language Bias in Multimodal Retrievers
+Language bias in multimodal retrieval refers to performance disparities where high-resource languages dominate rankings for semantically equivalent queries.
+osmulski2025miracl report that modern retrievers exhibit significant variation in NDCG scores [ndcg] across languages, indicating that their effectiveness depends on resources and script types.
+To quantify these disparities, fairness-aware metrics like exposure parity [ndcg] have been adapted, yet adewumi2024fairnessbiasmultimodalai emphasizes the lack of dedicated language-focused protocols.
+laosaengpha2025mitigatinglanguagebiascrosslingual introduced LBKL, a distributional measure of divergence for text modality bias.
+While technically extensible to multimodal retrieval, such metrics ignore retrieval rankings in the bias measurement.
+2.2 Cultural Benchmarks
+While recent benchmarks [liu-etal-2021-visually, romero2024cvqa, nayak-etal-2024-benchmarking, li2025ravenea] have made progress in evaluating cultural reasoning and knowledge, they primarily focus on culture-specific concepts or broad facets.
+In contrast, the assessment of bias discussed in Section 2.1 requires a basic-level, parallel corpus of universal concepts to strictly disentangle semantic relevance from cultural association in retrieval rankings.
+We provide a detailed comparison to other datasets in Appendix A.
+3 Methodology
 This section outlines the framework developed to investigate perspectival bias in multilingual, multimodal retrieval systems.
-We first state our guiding research questions and then explain the studies that address these research questions in Sections 3.1 and 3.2.
-Our investigation centers on two complementary perspectives, corresponding to different retrieval directions, as shown in Figure 2:
+We first state our guiding research questions and subsequently detail the experiments addressing them in Sections 3.1, 3.2.2, and 3.2.3.
+Our investigation is organized into three specific research questions as follows:
   1. RQ1
@@
 […]
-Figure 2: Overview of the study.
-First, in RQ1, we identify language prevalence in image-to-text retrieval by analyzing the language of the retrieved text and comparing it with high-resource languages, as well as medium- and low-resource languages.
-Second, in RQ2, we identify the association bias using self-preference behavior of the model by retrieving an image with three candidates: semantically relevant, culturally relevant, and a non-relevant candidate.
-3.1 RQ1: Image-to-Text Retrieval Study
-In this study, we assess the bias arising from linguistic prevalence by examining the discrepancy between an expected “fair” language distribution and the observed one.
-That is, the discrepancy should be zero if the linguistic prevalence has no effect on the retrieval results and increases as the results deviate from the ideal case.
-As discussed in Section 2.2, existing work lacks a dedicated metric to quantify such a discrepancy in multimodal retrieval.
-As the first step to closing this gap, we apply the Language Bias Kullback–Leibler (LBKL) Divergence proposed by Laosaengpha et al.
-[8], which measures the divergence between an expected language distribution and the observed distribution in a retrieved list.
-Given proportions of language A and B in the ground truth ($P_{A}(x),P_{B}(x)$) and in the retrieved set ($Q_{A}(x),Q_{B}(x)$), LBKL is given as:
+  3. RQ3
+[…]
+3.1 Image-to-Text Retrieval (RQ1)
+To assess prevalence bias, we measure the divergence between expected and observed language distributions.
+We adopt the Language Bias Kullback–Leibler (LBKL) divergence [laosaengpha2025mitigatinglanguagebiascrosslingual]:
 $\text{LBKL}=\frac{\sum\limits_{i=1}^{q}\left[P_{\text{A}}(x)\log\frac{P_{\text{A}}(x)}{Q_{\text{A}}(x)}+P_{\text{B}}(x)\log\frac{P_{\text{B}}(x)}{Q_{\text{B}}(x)}\right]}{q}$ (1)
-While LBKL can be applied to cross-modal retrieval, it is rank-agnostic: deviations at rank 1 are penalized equally to deviations at rank 100.
-This underestimates the harm in systems that concentrate resource-driven bias near the top ranks.
-For the next step, we extend LBKL by introducing the Discounted Language Bias Kullback-Leibler (DLBKL) divergence, which incorporates a logarithmic rank discount inspired by NDCG [18].
-We assign a weight $w(i)=1/\log_{2}(i+1)$ to each rank $i$.
-The rank-weighted proportion for a language $l$ is then:
+Since LBKL is rank-agnostic, we propose Discounted LBKL (DLBKL) to prioritize penalizing bias in higher ranks more than lower ranks.
+We apply a logarithmic discount $w(i)=1/\log_{2}(i+1)$ to the observed language proportion $Q^{\prime}_{l}(x)$:
 $Q^{\prime}_{l}(x)=\frac{\sum_{i=1}^{k}w(i)\cdot\mathbb{I}(\text{doc}_{i}\text{ is }l)}{\sum_{i=1}^{k}w(i)}$ (2)
-where $\mathbb{I}(\cdot)$ is the indicator function.
-DLBKL is calculated by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in the LBKL formula.
-As illustrated in Figure 3, DLBKL penalizes top-ranked disparities more heavily, aligning the metric with user exposure and better capturing the discrepancy between the ideal case and observed one in multimodal retrieval.
-Figure 3: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists where high-resource languages dominate the top ranks.
-3.2 RQ2: Text-to-Image Retrieval Study
-Figure 4: Overview of the XCM dataset creation process, designed to produce a benchmark with parallelism across semantics, cultures, and languages.
+DLBKL is computed by substituting $Q^{\prime}_{l}(x)$ for $Q_{l}(x)$ in equation 1.
+This formulation can be generalized to support cases where the bias assessment involves more than two languages by measuring divergence over the full distribution.
+It penalizes models where high-resource languages disproportionately occupy top ranks relative to the prior $P$.
+Crucially, in parallel datasets where semantic relevance is uniform across languages, this formulation isolates prevalence bias without compromising retrieval accuracy (see Figure 2 and Appendix D).
+Figure 2: Illustration of how DLBKL, unlike the rank-agnostic LBKL, assigns a higher bias score to lists in which high-resource languages dominate the top ranks, under the setup that the prior P is a uniform distribution
+3.2 Text-to-Image Retrieval (RQ2 & RQ3)
 To quantify the degree to which models prioritize cultural association over semantic fidelity, a phenomenon we illustrate in Figure 1(b), a benchmark with a parallel structure in its cultural dimension is necessary.
 To the best of our knowledge, no such benchmark exists, so we make two primary contributions.
-First, we construct and introduce the Cross-Cultural Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose.
+First, we construct and introduce the Cross-Cultural, Cross-Modal, Cross-lingual Multimodal (3XCM) benchmark, a novel dataset designed specifically for this purpose.
 Second, we propose the Self-Preference Cultural Bias Score (SP), a new metric for explicitly measuring this form of bias.
 3.2.1 The 3XCM Dataset Benchmark
-To evaluate association bias, we constructed the 3XCM benchmark [^1] Research release only (CC BY-NC-SA 4.0). Ethical review required for production use.
-Available at: https://huggingface.co/datasets/Chula-AI/association_bias_benchmark.
-. The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
+Figure 3: Overview of the XCM dataset creation process, designed to produce a benchmark with parallelism across semantics, cultures, and languages.
+To evaluate association bias, we constructed the 3XCM benchmark, a challenge set designed to elicit culture-language association bias.
+This evaluation requires strict cultural parallelism at the concept level (e.g., "school" or "wedding" across all cultures), a structural feature that existing datasets [liu-etal-2021-visually, romero2024cvqa, nayak-etal-2024-benchmarking] do not provide.
+[^1] Research release only (CC BY-NC-SA 4.0).
+Ethical review required for production use.
+The process involved two primary stages: (i) gathering a corpus of culturally diverse images and (ii) structuring these images into a triplet-based evaluation set.
 The image gathering stage, summarized in Figure 3, consisted of three steps:
@@
 Each concept is an abstract, semantic category that uses shared properties to group a broad, culturally-inclusive range of entities.
-The prompt for generating concepts can be found in Appendix A.
-  • Concept De-duplication: We use BGE-M3 [21] to de-duplicate concepts based on similarity with a threshold of 0.92.
-  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [22] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
+The prompt for generating concepts can be found in Appendix H.
+  • Concept De-duplication: We use BGE-M3 [chen2024bgem3embeddingmultilingualmultifunctionality] to de-duplicate concepts based on similarity with a threshold of 0.92.
+  • Image Collection: For each concept and a set of 16 diverse countries, we used the DuckDuckGo image search API [deepanprabhu:duckduckgo] to retrieve the top 10 images using queries in both English (e.g., "train Japan") and the local native language.
   • Image De-duplication: To ensure visual diversity, we performed two-stage de-duplication within each concept.
 First, near-exact duplicates were removed automatically using an embedding model.
-Subsequently, three human annotators, following the guidelines in Appendix F, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
-Leveraging the collected cultural images, we introduce a novel evaluation paradigm that employs a forced-choice task.
-This setup is designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association.
-As illustrated in Figure 5, for a given query (e.g., "food" in Thai), the model is presented with a triplet of image candidates: (i) Semantically Relevant: same concept, different culture (e.g., Nigerian food); (ii) Culturally Relevant: different concept, same culture (e.g., Thai traditional dance); and (iii) Non-Relevant: different concept and culture (e.g., Japanese gas station).
-Figure 5: Illustration of association bias evaluation. A Thai text query for “food” is evaluated against three candidates designed to isolate semantic faithfulness vs. cultural relevance.
-The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix J and N respectively.
-3.2.2 Self-Preference Cultural Bias Score (SP)
-With the constructed dataset, we can now measure the discrepancy between the ideal case and the observed one, where bias arising from cultural association may intervene.
+Subsequently, three human annotators, following the guidelines in Appendix J, used a custom tool to manually filter out remaining images that depicted the same scene or object without meaningful variation in viewpoint or time of day.
+The final dataset contains 11,724 entries distributed across 138 concepts. Further statistics and samples are provided in Appendix K and L respectively.
+3.2.2 RQ2: Effect of Implicit Association Bias
+To evaluate implicit association bias, we set up a forced-choice task designed to disambiguate between the model’s reliance on semantic understanding (the concept) and its preference for cultural association.
+As illustrated in Figure 4(a), for a given query (e.g., "train" in Thai), the model is presented with a triplet of image candidates representing three answer categories: (i) Correct: The model successfully identifies an image with a semantically relevant object (e.g., a train of any culture).
+(ii) Language-biased: The model prioritizes culture-language association over semantic relevance (e.g., a Thai dance).
+(iii) Totally irrelevant: The model retrieves an image without any semantic or cultural relevance (e.g., the Eiffel Tower).
+Figure 4: Illustration of association bias evaluation.
+(a) RQ2 evaluates a Thai query for “train” against three candidates to distinguish semantic faithfulness from cultural relevance.
+(b) RQ3 uses the query “train of Japan” to measure the tension between explicit cultural descriptors and implicit language-cultural associations.
+Note that we deliberately exclude images with relevance in terms of object semantics and query language to expose the prioritization of relevance.
+With the constructed dataset, we can now measure the discrepancy between ideal retrieval outcomes and observed results, where bias arising from cultural association may intervene.
 Ideally, the discrepancy should be zero when image retrieval depends solely on semantic relevance, and it should increase as the model’s preference tends towards images associated with the culture of the query, rather than semantic accuracy.
 To quantify the discrepancy, we propose a metric called the self-preference cultural bias score (SP), which can be computed as follows:
-$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{sem},i},S_{\text{cul},i},S_{\text{non},i})\right)$ (3)
-$\text{SP}=\frac{M_{\text{cul}}}{M_{\text{sem}}}$ (4)
+$M_{k}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(S_{k,i}=\max(S_{\text{cr},i},S_{\text{lb},i},S_{\text{ti},i})\right)$ (3)
+$\text{SP}=\frac{M_{\text{lb}}}{M_{\text{cr}}}$ (4)
 where $M_{k}$ is the proportion of times a candidate of type $k$ receives the highest similarity score across $N$ total trials.
-The candidate type $k$ can be semantically relevant (sem), culturally relevant (cul), or non-relevant (non).
+The candidate type $k$ can be Correct (cr), Language-biased (lb), or Totally Irrelevant (ti).
 The similarity score for candidate type $k$ in trial $i$ is denoted by $S_{k,i}$.
 The indicator function $\mathbb{I}(\cdot)$ is 1 if the condition is true and 0 otherwise.
-The SP score (Eq.
-4) is then the ratio of cultural wins ($M_{\text{cul}}$) to semantic wins ($M_{\text{sem}}$).
+The SP score (equation 4) is then the ratio of language-biased wins ($M_{\text{lb}}$) to correct answer wins ($M_{\text{cr}}$).
 In this way, a higher SP score indicates stronger cultural self-preference over semantic faithfulness and thus a greater extent of association bias.
+3.2.3 RQ3: Effect of Cultural Descriptors
+To investigate the model’s adherence to explicit instructions against language-driven bias, we extend the forced-choice framework to include specific cultural cues.
+We construct queries using an “Object & Cultural Descriptor” format (e.g., “Train of Japan” formulated in Thai) to gauge the “tugging effect” between the explicit country mention and the implicit language-cultural association.
+As illustrated in Figure 4(b), we expand the answer category to six to isolate specific retrieval failures: (i) Correct: The model successfully identifies an image with a semantically relevant object and cultural descriptor(e.g., a Shinkansen).
+(ii) Object Relevant, Language-biased: The model successfully identifies an image with a semantically relevant object, but is distracted by the culture associated with the query language (e.g., a Thai train); (iii) Object Relevant: The model successfully identifies an image with a semantically relevant object, but the culture is irrelevant to the query language and the cultural descriptor (e.g., British Train).
+(iv) Cultural Descriptor Relevant: The model successfully identifies an image associated with the cultural descriptor, but misses the object’s semantics (e.g., Sushi).
+(v) Language-biased: The model fails to identify an image with a semantically relevant object but matches the query language’s culture (e.g., Thai dance).
+(vi) Totally Irrelevant: The model captures neither the relevant object nor the culture associated with the descriptor or language.
+(e.g., Eiffel Tower).
+To quantify the "tugging effect" between implicit language bias and explicit instructions, we calculate the Similarity Drift ($\Delta\text{sim}$).
+Given a base concept query $q_{C}$ and a culturally-augmented query $q_{C+CD}$, the drift for an image $I$ is defined as:
+$\Delta\text{sim}(I)=S_{q_{C+CD},I}-S_{q_{C},I}$ (5)
+A higher $\Delta\text{sim}$ for the Correct category relative to other categories with mismatch culture indicates a model’s ability to prioritize explicit descriptors over inherent culture-language associations.
 4 Experimental Setup
-To answer our research questions, we conducted two main experiments.
-For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [23].
-The dataset offers a balanced multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities.
+To answer our research questions, we conducted three main studies.
+For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [ThapliyalCrossmodal2022].
+The dataset offers a parallel multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities.
+With this dataset, we compute Language-wise DLBKL by defining the prior $P(x)$ as a uniform distribution (see Appendix E)".
 We evaluate models using Accuracy@5, NDCG@10, LBKL@10, and our proposed DLBKL@10.
-For RQ2, we performed text-to-image retrieval on our newly created XCM benchmark, evaluating models using our proposed SP score.
-For both RQ1 and RQ2, we selected a representative suite of models spanning three distinct architectural paradigms, as shown in Table 1:
-  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data.
-We include the original CLIP-L/14 as a powerful baseline, and Chinese-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
-  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space.
-We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
-  • MLLM-Based Retrieval Embedders: This modern paradigm adapts large, pre-trained Multimodal Language Models for retrieval.
-We evaluate several state-of-the-art models, including the ColQwen series (v0.2, 3b-M, 7b-M), GME models (Qwen2-2B, Qwen2-7B), and Jina-E-v4.
-Full model identifiers are available in Appendix H.
+For RQ2, we performed text-to-image retrieval on our newly created 3XCM benchmark, evaluating models using our proposed SP score.
+For RQ3, we extend text-to-image retrieval of RQ2 by using the extended version of the 3XCM benchmark with six candidates and analyze the similarity drift.
+We selected a representative suite of models spanning three distinct architectural paradigms: Vision-Language Contrastive Models: foundational models trained with a contrastive objective mainly on English data, Cross-lingual Alignment Models: models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space for multilingual capability, and MLLM-Based Retrieval Embedders: utilize MLLM as the backbone and finetune for a contrastive objective for retrieval.
+We refer readers to Appendix C for detailed architectural descriptions and specific model configurations.
 5 Experimental Results
@@
 5.1 Image-to-Text Evaluation (RQ1)
-All models exhibit some degree of linguistic prevalence bias.
-For most MLLM-based models (Jina, ColQwen, GME), the DLBKL score is higher than the LBKL score, as shown in Table 1.This confirms that bias is more pronounced at the top of the ranked list, as these models tend to rank results from medium-to-high resource languages.
-Results for additional ranks and an example of retrieval result can be found in Appendix D.
+All models exhibit some degree of linguistic prevalence bias, as shown in Table 1.
+This bias is most pronounced in the top-ranked results, as models tend to prioritize high-resource languages over others in the early ranks.
+Appendix G provides visualizations revealing the dominance of high-resource languages in the top ranks and the overall disparity in retrieval frequency between language resource tiers.
+Results for additional ranks and an example of a retrieval result can be found in Appendix F.
 Model Acc LBKL DLBKL NDCG
@@
 Vision-Language Contrastive Models
-CLIP-L/14 0.509 5.673 5.684 0.290
-Chinese-CLIP-L/14 0.355 5.046 5.055 0.207
+CLIP-L/14 0.509 15.394 15.398 0.290
+CN-CLIP-L/14 0.355 15.287 15.291 0.207
 Cross-lingual Alignment Models
-XLM-R-L/14 0.924 0.320 0.333 0.736
-XLM-R-B/16plus 0.968 0.110 0.125 0.791
+XLM-R-L/14 0.924 12.775 12.793 0.736
+XLM-R-B/16plus 0.968 12.651 12.669 0.791
 MLLM-Based Retrieval Embedders
-ColQwen2.5-3b-M 0.894 0.792 0.817 0.605
-ColQwen2.5-7b-M 0.926 0.821 0.849 0.665
-ColQwen2.5-v0.2 0.754 3.834 3.867 0.481
-GME-Qwen2-2B 0.967 3.121 3.174 0.717
-GME-Qwen2-7B 0.979 1.371 1.420 0.770
-Jina-E-v4 0.972 0.915 0.951 0.775
-Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL. Explicit alignment models (XLM-R) show substantially lower bias.
-This phenomenon is visualized in Figure 6, which shows a clear dominance of high-resource languages in the top ranks.
-This further illustrates the overall disparity in retrieval frequency between language resource tiers as shown in Figure 7.
-Figure 6: Distribution of language groups across retrieval ranks. High-resource languages (blue) dominate the top ranks, a bias captured by DLBKL.
-Figure 7: Histogram of retrieved language frequencies. MLLM-based models disproportionately retrieve texts from medium-high resource languages (blue) over low-resource ones (orange).
+ColQwen2.5-3b-M 0.894 13.654 13.667 0.605
+ColQwen2.5-7b-M 0.926 13.439 13.454 0.665
+ColQwen2.5-v0.2 0.754 14.228 14.241 0.481
+GME-Qwen2-2B 0.967 13.627 13.644 0.717
+GME-Qwen2-7B 0.979 13.378 13.397 0.770
+Jina-E-v4 0.972 13.008 13.025 0.775
+Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL, calculated in a language-wise manner. Explicit alignment models (XLM-R) show substantially lower bias.
 Crucially, the explicit alignment models (XLM-R series) achieve the lowest bias scores by a significant margin, with XLM-R-B/16plus demonstrating near-zero linguistic prevalence bias according to both metrics, while maintaining high retrieval accuracy.
@@
 5.2 Text-to-Image Evaluation (RQ2)
-Using the proposed XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models.
-In this evaluation, the semantic win rate ($M_{\text{sem}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias.
-We observe that the baseline CLIP and Chinese-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
+Using the proposed 3XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models.
+In this evaluation, the semantic win rate ($M_{\text{cr}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias.
+We observe that the baseline CLIP and CN-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
 […]
@@
 CLIP-L/14 51.24% 40.78% 7.98% 0.80
-Chinese-CLIP-L/14 56.39% 31.65% 11.95% 0.56
+CN-CLIP-L/14 56.39% 31.65% 11.95% 0.56
 Cross-lingual Alignment Models
@@
 XLM-R-B/16plus 87.54% 6.23% 6.24% 0.07
-LLM-Based Retrieval Embedders
+MLLM-Based Retrieval Embedders
 GME-Qwen2-2B 83.34% 11.64% 5.02% 0.14
@@
 The CLIP-L/14 model, lacking a robust understanding of non-Latin scripts, defaults to matching cultural origin as a retrieval heuristic.
-Training on a large Chinese dataset (Chinese-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters.
+Training on a large Chinese dataset (CN-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters.
 However, this is a shallow fix that fails to generalize to other non-Latin scripts.
@@
 This challenge with low-resource languages persists even in more advanced architectures.
-For instance, MLLM-based models (Jina-E-v4) employ a LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages.
+For instance, MLLM-based models employ an LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages.
 For the vision component, a Vision-Language Model (VLM) is used as the image encoder to improve contextual awareness.
@@
 Figure 5: Association bias evaluation across four models reveals the limitations of monolingual training.
-The baseline CLIP (a) shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in Chinese-CLIP (b).
-In contrast, cross-lingual models like XLM-R (c) and particularly Jina-E-v4 (d) prove far more effective at mitigating this bias and maintaining high semantic relevance across diverse countries.
-Figure 9: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) Chinese-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
-This behavior is clearly visualized in the UMAP [24] projections of the text embeddings as shown in Figure 9.
+The baseline CLIP shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in CN-CLIP.
+In contrast, cross-lingual models like XLM-R and particularly Jina-E-v4 prove far more effective at overriding this bias and maintaining high semantic relevance across diverse countries.
+Figure 6: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) CN-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
+This behavior is clearly visualized in the UMAP [McInnes2018] projections of the text embeddings as shown in Figure 6.
 The baseline CLIP-L/14 model exhibits a fractured embedding space, with non-Latin languages forming distinct clusters far from the main Latin-script cluster.
 This demonstrates a lack of shared semantic understanding.
-In the Chinese-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated.
+In the CN-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated.
 In contrast, the explicit alignment model, XLM-R-L/14, successfully unifies the embedding space into a single, language-agnostic cluster, demonstrating a truly shared semantic representation across scripts.
@@
 This suggests a more flexible alignment that may capture finer semantic nuances between languages.
-To validate these visual findings numerically, we calculated the silhouette score [25] for each language’s text embeddings.
-This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix I.
+To validate these visual findings numerically, we calculated the silhouette score [shahapure2020cluster] for each language’s text embeddings.
+This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix N.
 This quantitatively reinforces that poor semantic understanding in the text encoder (as visualized by the disparate UMAP clusters) is a key driver of higher cultural association bias.
@@
 However, neither paradigm consistently outperforms the other on this specific task.
+Model behavior remains consistent regardless of prompt phrasing, such as verbose descriptions (e.g., “a picture of a train for my homework”) or culture-agnostic instructions (e.g., “focus on semantics instead of textual language”) (Appendices R.1 and S).
+5.3 Cultural Descriptor Text-to-Image Evaluation (RQ3)
+To investigate whether explicit instruction can override implicit association bias, we evaluate the models on the Extended 3XCM benchmark using queries with explicit Cultural Descriptors (CD) (e.g., “Train of Japan” in Thai).
+Adherence to Cultural Descriptors.
+Quantitative results demonstrate that most models are responsive to explicit cultural cues.
+As shown in Table 3, the correct answer (matching both object and cultural descriptor) is the most frequently selected candidate across all high-performing models (XLM series, MLLM-based).
+This suggests that explicit cultural descriptors (e.g., the word “Japan”) exert a stronger influence on the retrieval ranking than the implicit bias of the query language (e.g., Thai script), provided the model has sufficient multilingual capacity.
+Figure 7: Comparative association bias across CLIP-L/14, CN-CLIP-L/14, Jina-embeddings-v4, and XLM-R-L/14 models in 16 countries.
+Stacked bars show win percentages for six categories, ranging from Correct to Total Irrelevance, highlighting performance variance across diverse cultural contexts.
+Model $M_{\text{cr}}$ $M_{\text{orlb}}$ $M_{\text{or}}$ $M_{\text{cdr}}$ $M_{\text{lb}}$ $M_{\text{ti}}$
+Vision-Language Contrastive Models
+CLIP-L/14 43.05 29.83 5.30 8.54 12.56 0.72
+CN-CLIP-L/14 41.97 24.54 9.81 10.04 11.21 2.42
+Cross-lingual Alignment Models
+XLM-R-L/14 77.83 7.21 6.85 7.55 0.26 0.30
+XLM-R-B/16plus 72.98 9.88 9.70 6.57 0.40 0.48
+MLLM-Based Retrieval Models
+GME-Qwen2-2B 60.01 18.38 11.49 8.05 1.56 0.50
+GME-Qwen2-7B 59.75 20.84 10.79 6.81 1.38 0.43
+ColQwen2.5-v0.2 54.77 20.15 12.57 9.56 2.13 0.82
+ColQwen2.5-3B-M 58.73 20.71 9.26 9.37 1.31 0.62
+ColQwen2.5-7B-M 63.19 17.70 7.82 9.80 1.11 0.38
+Jina-E-v4 57.76 19.59 13.62 7.14 1.20 0.68
+Table 3: Association bias results on the 3XCM benchmark (6 categories).
+$M_{\text{cr}}$, $M_{\text{orlb}}$, $M_{\text{or}}$, $M_{\text{cdr}}$, $M_{\text{lb}}$, and $M_{\text{ti}}$ represent Correct, Object Relevant with Language Bias, Object Relevant, Cultural Descriptor Relevant, Language-biased, and Totally Irrelevant, respectively.
+Failure Mode Analysis.
+Let us examine two failure modes presented by the error distribution in Figure 7 First, when the model successfully retrieves the correct semantic object but fails to adhere to the explicit cultural descriptor (i.e., "Semantic Success, Cultural Failure"), we observe that the selection rate of the Object Relevant Language-biased category (which matches the object and the query language’s culture, represented by purple color) is higher than that of the Object Relevant category (which matches the object with an unrelated culture, represented by blue color).
+This difference is significant ($p$-value < 0.05, Chi-squared test) for most models except the XLM family (see Appendix P.2).
+The trend is most pronounced in non-Latin and low-resource languages, suggesting that when the explicit cultural constraint is violated, the model falls back on implicit language-driven associations.
+Second, when the model fails to retrieve the correct semantic object (i.e., "Semantic Failure"), the model is more likely to select the Cultural Descriptor Relevant category (which matches the explicit cultural descriptor only, represented by yellow color) than the Language-biased category (which matches the query language’s culture only, represented by pink color).
+This difference is significant ($p$-value < 0.05, Chi-squared test) for most models except the Vision-Language Contrastive group (CLIP and CN-CLIP), which lack robust non-Latin script understanding (see Appendix P.2).
+This indicates that when the model fails to ground the visual object, it prioritizes the explicit cultural token (e.g., the country name) over the implicit association of the script.
+(See Appendices P and R.2).
+Similarity Drift Analysis.
+The Similarity Drift ($\Delta\text{sim}$) metric measures the change in cosine similarity for a target image when a cultural descriptor is added to the query.
+On average, adding a cultural descriptor successfully shifts the similarity distribution in the positive direction (toward the Correct category) across most models and languages.
+This confirms that the models generally recognize and attend to the explicit instruction.
+However, a notable exceptions emerge.
+For the baseline CLIP model, queries in non-Latin scripts exhibit near-zero drift, indicating that the text encoder fails to process the cultural descriptor in these scripts, rendering the explicit instruction ineffective (see Appendix Q.1).
 6 Discussion
-Our evaluation framework distinguishes between two perspectival biases: prevalence bias, driven by data imbalance, and association bias, arising from learned cultural correlations.
-Our findings show these are distinct challenges.
-Explicit cross-lingual alignment, used by the XLM-R models, is a highly effective strategy, achieving the lowest scores for both prevalence bias (DLBKL) and association bias (SP) by directly enforcing a shared semantic space.
-While modern MLLMs like Jina-E-v4 also perform well against association bias, the persistence of these issues across all models points to a deeper, unresolved problem: the entanglement of semantic concepts with linguistic and cultural artifacts in the model’s embedding space.
-The path forward, therefore, requires a fundamental shift in training strategy.
-Future work must prioritize training objectives that actively enforce non-association by creating a truly global semantic space.
-This means designing models to map a semantic query, regardless of its language or cultural origin, to all conceptually relevant images, irrespective of their geographical context.
-For example, the new method must include a curation process to avoid cross-cultural false negative images being pushed away from their corresponding queries, and utilize data augmentation to help ensure language-agnostic property.
+Perspectival Biases and the Power of Explicit Alignment.
+Our investigation establishes that perspectival bias manifests in two correlated forms: prevalence bias (RQ1) from distributional imbalances in training data, favoring high-resource languages, and association bias (RQ2) from spurious correlations between language scripts and cultural visual features.
+While both arise from training data and correlate empirically (high DLBKL models tend toward high SP), explicit cross-lingual alignment effectively reduces both, as seen in XLM-R models’ low bias scores and high accuracy.
+This alignment not only mitigates prevalence and association biases but also enhances the model’s ability to follow explicit cultural descriptors (CDs).
+Guidance via Cultural Descriptors (CDs) and the Tugging Effect.
+CDs provide guidance to steer culture.
+The positive Similarity Drift ($\Delta\text{sim}$) in RQ3 shows models can adhere to cultural descriptors, steering retrieval toward target cultures in most cases despite the tugging effect of query-language associations.
+However, failure mode analysis reveals the tugging effect: when matching the object but not the CD, models revert to culture-language association over neutral alternatives.
+This indicates association bias as a latent prior suppressed but not eliminated by guidance.
+Modern MLLM-based models still exhibit issues with both biases and inconsistent compliance with CD guidance, especially for low-resource or non-Latin languages, where drifts are erratic or near-zero without robust multilingual support.
 7 Conclusion
-In this work, we introduce a framework that distinguishes between two forms of perspectival bias in multimodal retrieval, prevalence bias and association bias, reflecting distinct ways in which a model’s prior shapes its behavior.
-This conceptual framing is operationalized through our proposed metrics and datasets: DLBKL, which measures rank-aware language prevalence bias, and XCM, which quantifies association bias through cross-cultural image retrieval.
-Together, these tools enable systematic evaluation of how multimodal large language models inherit and express perspectival biases across languages and cultures.
-In image-to-text retrieval, prevalence bias arises when a model favors texts from high-resource languages.
-This problem is addressed by anchoring other languages to the prevalent ones through cross-lingual alignment.
-In text-to-image retrieval, association bias arises when a model favors images that are culturally associated with the query language rather than semantically faithful to their content.
-Such bias cannot be resolved through traditional cross-lingual alignment or by merely exposing the model to a wider range of cultural content during training.
-Ultimately, our findings call for a more principled approach: one that directly mitigates localized spurious association as a core design principle for models that are not only multilingual but also perform consistently across languages and cultures.
+We distinguish prevalence bias, a distributional skew, from association bias, a semantic entanglement between language and culture.
+Even with explicit cultural descriptors, models exhibit a systematic “tugging effect” to language-culture association bias when semantic grounding succeeds but cultural disambiguation fails.
+These findings demonstrate that achieving cross-cultural equity requires more than naive scaling.
+Future work must prioritize training paradigms that actively de-correlate language from cultural features, enforce cultural constraints for steerability, or explicit cross-lingual alignment.
+Only by compiling these can we achieve a retrieval system that truly operates across global contexts with cultural controllability.
 8 Limitations
@@
 It therefore cannot distinguish between retrieving irrelevant documents and over-representing a language with relevant ones.
-Second, the XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures.
-The benchmark’s coarse-grained semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity.
-Acknowledgments
-This research has received funding support from the National Science, Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B13F680099] and Chulalongkorn University Artificial Intelligence Institute.
-Appendix A Culturally Relevant Concept Identification
-To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection.
-The prompt used in this process is shown in Figure 11 to identify all labels associated with an image.
-Figure 10.
-list 100 concepts that unique and vary in these country including China, India, Japan, saudi arabia, France, German, Brazil,
-Kenya, Thailand, USA
-like this
-{"food":{"China":"Mala Xiang Guo", ..., "Thailand":"Padthai", "USA":"hamburger"}, "costume":{"China":"Hanfu", ...,
-"Thailand":"Sabai", "USA":"cowboy"}}
-Figure 10: The prompt given to Gemini to generate unique country-specific image concepts.
-Appendix B Multi-Concept Detection in Cultural Images
-To establish the self-preference cultural bias, the culturally relevant and non-relevant candidate images must not share concepts with the text label.
-We utilized Gemini2 with the following prompt in Figure 11 to identify all labels associated with an image.
-Please classify the following image by assigning them to one or more of the following cultural
-categories:
-   {category}.
-**Comprehensive Output Format (JSON):**
-output as JSON for example:
-  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
-in the categories please order by priority (high to low).
-Figure 11: The prompt given to Gemini for multi-label image classification, where the {category} placeholder is dynamically populated with the full list of categories.
-Appendix C Language Resources
-To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [26] as an approximation.
-Table 3 and 4 presents the resulting language composition for RQ1 and RQ2, consequently.
-Language Type Language ID Distribution (%)
+Second, the 3XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures.
+The benchmark’s universal basic-level semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity, as its usage is intended to be used as a challenging, controlled environment used to uncover biases.
+Appendix A Comparison with Other Datasets
+We compare 3XCM with existing VQA benchmarks in Table LABEL:tab:benchmark_comparison. The 3XCM dataset was developed to address the lack of cross-cultural parallelism in current resources.
+Feature MaRVL [liu-etal-2021-visually] CulturalVQA [nayak-etal-2024-benchmarking] CVQA [romero2024cvqa]
+Primary Task Visual Reasoning Visual QA Visual QA
+Languages 5 11 31
+Granularity Specific indigenous Broad facets Broad categories
+Parallelism Low (cultural-specific) Low (cultural-specific) Low (region-specific)
+Bias Target Reasoning Failure Knowledge Gaps Knowledge Gaps
+(a) Comparison of established cultural VQA benchmarks.
+Feature RAVENEA [li2025ravenea] 3XCM (Ours)
+Primary Task VQA & Image Cap. Cross-Modal Retrieval
+Languages 8 (Countries) 16
+Granularity Varies (VQA:cultural-specific, IC: too board) Universal basic-level
+Parallelism Low (Mixed) High (Cross-cultural)
+Bias Target Cultural Nuances Association Bias
+(b) Comparison with recent multi-task and cross-modal benchmarks.
+Table 4: Comprehensive Benchmark Comparison.
+Appendix B Language Resources
+To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [CommonCrawlLanguages] as an approximation.
+Table 5 and 6 present the resulting language composition for RQ1 and RQ2, respectively.
+Type Language ID Distribution
 High English en 43.9499
@@
 Table 5: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ1
-Name Full Name Language Resources
-USA United States of America English 43.950
-UK United Kingdom
-AUS Australia
-GER Germany German 5.569
-CHN China Chinese 4.878
-JPN Japan Japanese 4.915
-ESP Spain Spanish 4.542
-ARG Argentina
-FRA France French 4.327
-PRT Portugal Portuguese 2.337
-BRA Brazil
-SAU Saudi Arabia Arabic 0.672
-THA Thailand Thai 0.427
-IND India Hindi 0.200
-KEN Kenya Swahili 0.010
-NGA Nigeria Yoruba 0.001
+Code Country Language Resources
+US America English 43.950
+GB Great Britain
+AU Australia
+DE Germany German 5.569
+CN China Chinese 4.878
+JP Japan Japanese 4.915
+ES Spain Spanish 4.542
+AR Argentina
+FR France French 4.327
+PT Portugal Portuguese 2.337
+BR Brazil
+SA Saudi Arabia Arabic 0.672
+TH Thailand Thai 0.427
+IN India Hindi 0.200
+KE Kenya Swahili 0.010
+NG Nigeria Yoruba 0.001
 Table 6: Composition of Language Resources in the CommonCrawl Dataset ($C$-MAIN-2025-18) for the language experimented in RQ2
-Appendix D Example of Result from Image to Text Retrieval
-To further elaborate the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 12.
-We also provide result of LBKL and DLBKL score at other rank in Table 5.
-Figure 12: An Example of Result from Image to Text Retrieval
+Appendix C Evaluated Models and Details
+In this work, we assess three distinct architectural paradigms for multimodal retrieval. The specific characteristics of these families are as follows:
+  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data.
+We include the original CLIP-L/14 as a powerful baseline, and CN-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
+  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space for multilingual capability.
+We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
+  • MLLM-Based Retrieval Embedders: This modern paradigm adapts large, pre-trained Multimodal Language Models for retrieval.
+We evaluate several state-of-the-art models, including the ColQwen series (v0.2, 3b-M, 7b-M), GME models (Qwen2-2B, Qwen2-7B), and Jina-E-v4.
+Table 7 provides the precise aliases, full model names, parameter counts, and citations used throughout the paper.
+Alias Used in Paper Full Model Name Parameter
+Vision-Language Contrastive Models
+   CLIP-L/14 clip-vit-large-patch141 427.6M
+   CN-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
+Cross-lingual Alignment Models
+   XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
+   XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
+MLLM-Based Retrieval Embedders
+   ColQwen2.5-v0.2 ColQwen2.5-v0.23 3814.8M
+   ColQwen2.5-3B-M ColQwen2.5-3b-multilingual-v1.03 3994.6M
+   ColQwen2.5-7B-M ColQwen2.5-7b-multilingual-v1.03 8071.1M
+   GME-Qwen2-2B gme-Qwen2-VL-2B-Instruct4 2209.0M
+   GME-Qwen2-7B gme-Qwen2-VL-7B-Instruct4 7070.6M
+   Jina-E-v4 jina-embeddings-v45 3934.7M
+The models are based on the following works: 1) DBLP:journals/corr/abs-2103-00020 for CLIP-L/14; 2) chinese-clip for CN-CLIP-L/14; 3) faysse2025colpaliefficientdocumentretrieval for ColQwen2 models; 4) zhang2025gmeimprovinguniversalmultimodal for GME-Qwen2 models; 5) günther2025jinaembeddingsv4universalembeddingsmultimodal for Jina-E-v4; and 6) carlsson-etal-2022-cross for XLM-R-VL models.
+Table 7: Aliases Used in Paper and Corresponding Full Model Names and Parameters
+Appendix D Details on LBKL and DLBKL
+In Section 3.1, we utilize LBKL and DLBKL to measure prevalence bias. Here, we elaborate on the rationale behind these metrics.
+In an ideal, bias-free retrieval scenario, the distribution of languages in the retrieved list should match the expected ”fair” distribution (e.g., the ground truth distribution).
+LBKL measures the Kullback–Leibler divergence between these two distributions.
+In equation 1, $P(x)$ represents the proportion of a specific language in the ground truth, and $Q(x)$ represents the proportion in the retrieved set.
+However, standard LBKL treats a deviation at rank 1 the same as a deviation at rank 100.
+In multimodal retrieval, users rarely scroll far, making top-rank bias significantly more harmful.
+Our proposed DLBKL in equation 2 addresses this by applying a discount factor derived from NDCG [ndcg].
+By weighting the observed proportion $Q(x)$ by the inverse log of the rank, we ensure that high-resource language dominance at the top of the list results in a significantly higher divergence score than if that same dominance occurred lower in the list.
+Appendix E Formulation of Language-wise LBKL and DLBKL
+In this section, we detail the calculation of two bias metrics: Language-wise LBKL (Country-LBKL) and Language-wise DLBKL (Country-DLBKL).
+Both metrics quantify the divergence between the retrieved country distribution and an ideal uniform distribution.
+The key distinction lies in how the predicted distribution is constructed: Country-LBKL treats the retrieved documents as an unweighted set, while Country-DLBKL applies rank-based discounting to prioritize higher-ranked results.
+E.1 Ground Truth Distribution (Shared)
+Let $C=\{c_{1},c_{2},...,c_{N}\}$ be the set of target countries, where $N=|C|$. We define the ground truth distribution $P$ as:
+$P(c_{j})=\frac{1}{N},\quad\forall c_{j}\in C$ (6)
+E.2 Country-LBKL (Unweighted)
+Country-LBKL evaluates bias based on the frequency of countries in the top-$k$ retrieved documents $D=[d_{1},...,d_{k}]$, treating all rank positions equally.
+E.2.1 Predicted Distribution (Unweighted)
+The unweighted proportion $Q(c_{j})$ is calculated as:
+$Q(c_{j})=\frac{1}{k}\sum_{i=1}^{k}\mathbb{1}_{c_{j}}(d_{i})$ (7)
+where the indicator function $\mathbb{1}_{c_{j}}(d_{i})$ is:
+$\mathbb{1}_{c_{j}}(d_{i})=\begin{cases}1&\text{if }d_{i}\text{ is from }c_{j}\\ 0&\text{otherwise}\end{cases}$ (8)
+E.2.2 Score Calculation
+The Country-LBKL score is the KL divergence between $P$ and $Q$. We simplify the substitution of $P(c_{j})$ as follows:
+$\begin{split}\text{Country-LBKL}&=\sum_{c_{j}\in C}P(c_{j})\log\left(\frac{P(c_{j})}{Q(c_{j})}\right)\\ &=\frac{1}{N}\sum_{c_{j}\in C}\log\left(\frac{1}{N\cdot Q(c_{j})}\right)\end{split}$ (9)
+E.3 Country-DLBKL (Rank-Weighted)
+Country-DLBKL assigns higher importance to documents appearing earlier in the list using logarithmic discounting.
+E.3.1 Rank-Based Weighting
+For rank position $i$, the weight $w_{i}$ is:
+$w_{i}=\frac{1}{\log_{2}(i+1)}$ (10)
+The total weight normalization factor is:
+$W_{total}=\sum_{i=1}^{k}w_{i}=\sum_{i=1}^{k}\frac{1}{\log_{2}(i+1)}$ (11)
+E.3.2 Predicted Distribution (Weighted)
+The weighted distribution $Q^{\prime}(c_{j})$ is calculated as:
+$Q^{\prime}(c_{j})=\frac{\sum_{i=1}^{k}w_{i}\cdot\mathbb{1}_{c_{j}}(d_{i})}{W_{total}}$ (12)
+E.3.3 Score Calculation
+The Country-DLBKL score is the KL divergence between $P$ and the weighted distribution $Q^{\prime}$:
+$\begin{split}\text{Country-DLBKL}&=\sum_{c_{j}\in C}P(c_{j})\log\left(\frac{P(c_{j})}{Q^{\prime}(c_{j})}\right)\\ &=\frac{1}{N}\sum_{c_{j}\in C}\log\left(\frac{1}{N\cdot Q^{\prime}(c_{j})}\right)\end{split}$ (13)
+E.4 Implementation Note
+For both metrics, a small smoothing value is applied to $Q(c_{j})$ and $Q^{\prime}(c_{j})$ in implementation to ensure numerical stability when a country is not represented in the retrieval results.
+Appendix F Example of Prevalence Bias Evaluation
+To further elaborate on the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 8.
+We also provide the results of LBKL and DLBKL scores at other ranks in Table 8.
+Figure 8: Qualitative Comparison of Image-to-Text Retrieval in Crossmodal-3600.
+Two retrieval results share identical semantic proportions and LBKL scores.
+Despite this semantic equality, the DLBKL score differs, capturing the model’s implicit preference for specific linguistic groups in the ranking process.
 Model @5 @25 @50 @99
@@
 Vision-Language Contrastive Models
-CLIP-L/14 6.904 6.911 4.171 4.182 3.245 3.249 2.658 2.652
-Chinese-CLIP-L/14 6.220 6.229 3.793 3.798 3.172 3.168 2.582 2.570
+CLIP-L/14 15.846 15.848 14.706 14.711 14.158 14.162 13.629 13.631
+CN-CLIP-L/14 15.784 15.787 14.419 14.427 13.680 13.688 12.968 12.974
 Cross-lingual Alignment Models
-XLM-R-L/14 0.939 0.960 0.240 0.246 0.221 0.223 0.214 0.213
-XLM-R-B/16plus 0.692 0.713 0.043 0.049 0.030 0.033 0.019 0.019
+XLM-R-L/14 14.710 14.718 8.547 8.583 4.759 4.800 2.386 2.413
+XLM-R-B/16plus 14.654 14.662 8.248 8.285 4.345 4.388 2.007 2.036
 MLLM-Based Retrieval Embedders
-ColQwen2.5-3B-M 2.138 2.164 0.192 0.209 0.114 0.122 0.088 0.089
-ColQwen2.5-7B-M 2.373 2.397 0.212 0.232 0.127 0.138 0.083 0.089
-ColQwen2.5-v0.2 5.958 5.974 1.375 1.424 0.633 0.676 0.377 0.410
-GME-Qwen2-2B 6.014 6.037 0.739 0.804 0.254 0.306 0.140 0.175
-GME-Qwen2-7B 3.843 3.875 0.221 0.264 0.094 0.124 0.058 0.077
-Jina-E-v4 2.859 2.886 0.157 0.186 0.072 0.093 0.045 0.059
+ColQwen2.5-3B-M 15.010 15.016 11.231 11.253 9.326 9.347 7.839 7.846
+ColQwen2.5-7B-M 14.947 14.954 10.585 10.610 8.316 8.339 6.750 6.755
+ColQwen2.5-v0.2 15.340 15.346 12.176 12.199 10.561 10.586 9.315 9.331
+GME-Qwen2-2B 15.144 15.151 10.457 10.494 7.664 7.710 5.322 5.368
+GME-Qwen2-7B 15.035 15.042 9.766 9.805 6.503 6.552 4.024 4.068
+Jina-E-v4 14.781 14.789 9.320 9.354 6.079 6.120 3.734 3.769
 Table 8: Image-to-text retrieval bias on Crossmodal-3600, measured by LBKL and DLBKL at various retrieval depths (k).
-Appendix E Language and Rank Frequency Diagram
-To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix C, showing both their overall retrieval frequency as shown in Figure 13 and their frequency distribution across ranks as shown in Figure 14.
-Figure 13: A language frequency histogram of each language group for all model
-Figure 14: A frequency of language group at each rank for all model
-Appendix F Annotator Guideline
+Appendix G Language and Rank Frequency Diagram
+To illustrate bias in image-to-text retrieval, we present a visualization of language groups categorized by resource level as shown in Appendix B, showing both their overall retrieval frequency as shown in Figure 9 and their frequency distribution across ranks as shown in Figure 10.
+Figure 9: A language frequency histogram of each language group (see Appendix B) for all models
+Figure 10: A frequency of language group (see Appendix B) at each rank for all model
+Appendix H Culturally Relevant Concept Identification
+To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection.
+The prompt used in this process is shown in Figure 11.
+list 100 concepts that unique and vary in these country
+including China, India, Japan, saudi arabia, France,
+German, Brazil, Kenya, Thailand, USA
+like this
+{"food":{"China":"Mala Xiang Guo", ..., "Thailand":
+"Padthai", "USA":"hamburger"}, "costume":{"China":
+"Hanfu", ..., "Thailand":"Sabai", "USA":"cowboy"}}
+Figure 11: The prompt given to Gemini to generate unique country-specific image concepts.
+Appendix I Multi-Concept Detection in Images
+To establish the self-preference cultural bias, the culturally relevant and non-relevant candidate images must not share concepts with the text label.
+We utilized Gemini2 with the following prompt in Figure 12 to identify all labels associated with an image.
+Please classify the following image by assigning them to one or more of the following cultural categories: {category}. **Comprehensive Output Format (JSON):** output as JSON for example:
+  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
+in the categories please order by priority (high to low).
+Figure 12: The prompt given to Gemini for multi-label image classification, where the {category} placeholder is dynamically populated with the full list of categories.
+Appendix J Annotator Guideline
 The guideline we provide to the annotators is to remove duplicates across multiple views.
@@
 This 397 includes slight shifts, crops, or zooms of the same scene.
-Appendix G UMAP Analysis for Self-Preference Cultural Bias
-To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 15 and 16.
-The text embeddings cluster strongly by language, a proximity that supersedes semantic content.
-Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias.
-While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models.
-The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster.
-This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
-Figure 15: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CLIP-L/14 model applied to our dataset.
-Figure 16: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the Chinese-CLIP-L/14 model applied to our dataset.
-Appendix H Full Official Model Name
-In this paper, we use aliases for the model names for conciseness; the full names are provided in Table 6.
-Alias Used in Paper Full Model Name Parameter
-CLIP-L/14 clip-vit-large-patch141 427.6M
-Chinese-CLIP-L/14 Chinese-clip-vit-large-patch142 406.2M
-ColQwen2.5-v0.2 ColQwen2.5-v0.23 3814.8M
-ColQwen2.5-3B-M ColQwen2.5-3b-multilingual-v1.03 3994.6M
-ColQwen2.5-7B-M ColQwen2.5-7b-multilingual-v1.03 8071.1M
-GME-Qwen2-2B gme-Qwen2-VL-2B-Instruct4 2209.0M
-GME-Qwen2-7B gme-Qwen2-VL-7B-Instruct4 7070.6M
-Jina-E-v4 jina-embeddings-v45 3934.7M
-XLM-R-L/14 XLM-Roberta-Large-Vit-L-146 998.3M
-XLM-R-L/16plus XLM-Roberta-Large-Vit-B-16Plus6 768.9M
-The models are based on the following works: 1) Radford et al.
-[2] for CLIP-L/14; 2) Yang et al.
-[27] for Chinese-CLIP-L/14; 3) Faysse et al.
-[14] for ColQwen2 models; 4) Zhang et al.
-[15] for GME-Qwen2 models; 5) Günther et al.
-[20] for Jina-E-v4; and 6) Carlsson et al.
-[10] for XLM-R-VL models.
-Table 6: Aliases Used in Paper and Corresponding Full Model Names and Parameters
-Appendix I Correlation Analysis for Self-Preference Cultural Bias
-We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix G, impacts cross-modal retrieval.
-To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with self-preference cultural bias score (SP) cultural bias as shown in Table 8.
-We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores.
-For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 17.
-Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder as shown in Table 7.
-Figure 17: Comparison of Self-Preference Cultural Bias with Text and Image Silhouette
-Model TSC ISC
-CLIP-L/14 0.83 0.55
-Chinese-CLIP-L/14 -0.26 0.58
-XLM-R-VL-B/16 0.98 -0.27
-XLM-R-VL-L/14 0.94 0.05
-Jina-E-v4 0.86 0.16
-GME-Qwen2-2B 0.80 0.09
-GME-Qwen2-7B 0.64 0.07
-Average 0.68 0.18
-Table 7: This table presents a Pearson correlation analysis between model performance and bias.
-We measure the correlation between association and the quality of data clusters (via Silhouette Score) in both the text embedding space (TSC) and the image embedding space (ISC).
-Model Metrics Country
-USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
-CLIP-L/14 SP $\downarrow$ 0.01 0.02 0.02 0.76 2.63 1.94 0.24 0.19 0.34 0.39 0.51 10.71 8.09 15.88 2.04 2.27
-TS $\downarrow$ 0.05 0.05 0.05 0.02 0.16 -0.05 -0.01 -0.02 -0.02 -0.01 -0.01 0.25 0.23 0.23 0.03 0.13
-IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.04 0.01 0.01 0.00 0.00 0.01 0.05 0.03 0.03 0.02 0.03
-Chinese-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
-TS $\downarrow$ -0.05 -0.05 -0.05 0.00 0.13 -0.07 -0.03 -0.03 -0.03 -0.01 -0.01 -0.40 0.93 -0.37 0.00 0.03
-IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.00 0.00 0.00 0.04 0.05 0.04 0.02 0.01
-Jina-E-v4 SP $\downarrow$ 0.02 0.03 0.02 0.05 0.04 0.04 0.04 0.03 0.05 0.04 0.04 0.05 0.02 0.12 0.49 0.93
-TS $\downarrow$ -0.01 -0.01 -0.01 -0.02 0.01 0.00 -0.01 -0.02 -0.02 -0.02 -0.02 0.02 0.01 0.00 0.00 0.13
-IS $\downarrow$ -0.01 0.00 -0.01 -0.01 0.00 0.01 0.00 0.00 -0.01 -0.01 -0.01 0.02 0.00 0.01 0.01 0.00
-XLM-R-L/14 SP $\downarrow$ 0.05 0.04 0.07 0.05 0.04 0.07 0.08 0.08 0.09 0.08 0.07 0.05 0.04 0.04 0.11 0.61
-TS $\downarrow$ -0.18 -0.18 -0.18 -0.11 -0.07 -0.10 -0.16 -0.16 -0.16 -0.12 -0.12 -0.07 -0.06 -0.17 -0.14 0.44
-IS $\downarrow$ 0.01 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.01 0.01 0.01 0.06 0.04 0.03 0.02 0.03
-XLM-R-B/16plus SP $\downarrow$ 0.05 0.04 0.06 0.04 0.04 0.05 0.05 0.06 0.07 0.06 0.06 0.04 0.02 0.05 0.10 0.66
-TS $\downarrow$ -0.24 -0.24 -0.24 -0.21 -0.19 -0.19 -0.24 -0.23 -0.23 -0.23 -0.23 -0.19 -0.19 -0.23 -0.23 0.49
-IS $\downarrow$ 0.01 0.01 0.00 0.00 0.01 0.01 0.00 0.00 -0.01 0.00 -0.01 0.03 0.02 0.03 0.02 0.00
-GME-Qwen2-2B SP $\downarrow$ 0.02 0.03 0.02 0.10 0.04 0.04 0.05 0.05 0.08 0.08 0.09 0.14 0.10 0.24 1.24 2.02
-TS $\downarrow$ 0.02 0.02 0.02 0.04 0.02 0.00 0.04 0.01 0.01 0.03 0.03 0.01 0.03 0.00 0.03 0.15
-IS $\downarrow$ 0.00 0.00 0.01 0.00 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.02 0.01 0.01
-GME-Qwen2-7B SP $\downarrow$ 0.03 0.02 0.01 0.14 0.04 0.05 0.07 0.07 0.07 0.09 0.13 0.14 0.08 0.12 0.62 1.88
-TS $\downarrow$ 0.04 0.04 0.04 0.05 0.02 0.02 0.07 0.02 0.02 0.04 0.04 0.09 0.11 0.03 0.04 0.14
-IS $\downarrow$ 0.00 0.00 0.01 0.00 0.01 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.01 0.01 0.01
-Table 8: Cross-Country and Cross-Model Comparison of Language and Cultural Bias Metrics.
-This table presents the results for the Self-Preference Cultural Bias score (SP), Text Silhouette (TS), and Image Silhouette (IS) scores across various multimodal retrievers for a selection of countries.
-Appendix J Dataset Statistics
-The distribution of cultural concepts in the XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
-Country Number of Samples
+Appendix K Dataset Statistics
+The distribution of cultural concepts in the 3XCM dataset is shown in Table 10, with each concept being represented by approximately 85 images on average.
+Country Samples
 Argentina 771
@@
 USA 609
-Average 733
-Table 9: Dataset Statistics of XCM dataset per Country
+Average 732.75
+Table 9: Dataset Statistics of 3XCM per Country
 Concepts (A-G)
@@
 Table 10: Distribution of Concepts and Image Counts
+Appendix L 3XCM Dataset Benchmark Samples
+This research provides an association evaluation benchmark and image metadata. Examples of the benchmark for RQ2, RQ3, and image metadata are shown in Figure 13, 14, and 15 respectively.
+Figure 13: Benchmark examples from the 3XCM dataset featuring three retrieval candidates.
+Figure 14: Benchmark examples from the 3XCM dataset featuring six retrieval candidates.
+Figure 15: Metadata for image of 3XCM dataset benchmark
+Appendix M UMAP Analysis for Self-Preference Cultural Bias
+To visualize cultural bias, the UMAP projections of text and image embeddings of all models as shown in Figure 16 and 17.
+The text embeddings cluster strongly by language, a proximity that supersedes semantic content.
+Conversely, the image embeddings do not exhibit strong country-based clustering, suggesting lower cultural bias.
+While other models show a similar, albeit less severe, tendency for text embeddings to be more biased than image embeddings, this effect is diminished in modern models.
+The GME-Qwen2 and Jina-E-v4 models only cluster very low-resource languages (Swahili, Yoruba), and the XLM-R models demonstrate superior alignment, forming a single central cluster.
+This discrepancy challenges retrieval systems: a query’s text embedding is biased by its language, leading the system to favor images from the same cultural context over potentially more visually relevant content from others.
+Figure 16: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CLIP-L/14 model applied to our dataset.
+Figure 17: The UMAP visualizations of the caption embeddings (left) and image embeddings (right) from the CN-CLIP-L/14 model applied to our dataset.
+Appendix N Correlation Analysis for Self-Preference Cultural Bias
+We investigate how unimodal bias, which our UMAP analysis shows is more severe in the text modality as shown in Appendix M, impacts cross-modal retrieval.
+To quantify this, we use the Silhouette score and find that high scores in low-resource languages correlate with the self-preference cultural bias score (SP), as shown in Table 12.
+We confirm this relationship by calculating the Pearson correlation between SP and the Silhouette scores.
+For example, CLIP-L/14’s Text Silhouette score correlates strongly with SP (0.827), while its Image Silhouette correlation is only moderate (0.550), as shown in Figure 18.
+Across all tested models, the average correlations reveal that SP is predominantly driven by the text encoder, as shown in Table 11.
+Figure 18: Comparison of Self-Preference Cultural Bias with Text and Image Silhouette
+Model TSC ISC
+CLIP-L/14 0.83 0.55
+CN-CLIP-L/14 -0.26 0.58
+XLM-R-VL-B/16 0.98 -0.27
+XLM-R-VL-L/14 0.94 0.05
+Jina-E-v4 0.86 0.16
+GME-Qwen2-2B 0.80 0.09
+GME-Qwen2-7B 0.64 0.07
+Average 0.68 0.18
+Table 11: This table presents a Pearson correlation analysis between model performance and bias.
+We measure the correlation between association and the quality of data clusters (via Silhouette Score) in both the text embedding space (TSC) and the image embedding space (ISC).
 Model Metrics Country
-USA UK AUS GER CHN JPN FRA ESP ARG PRT BRA SAU THA IND KEN NGA
-CLIP-L/14 $M_{\text{sem}}(\%)\uparrow$ 95.73 94.57 94.73 52.42 25.17 31.07 75.53 78.00 69.65 65.78 61.33 7.75 10.48 5.56 27.83 24.19
-$M_{\text{cul}}(\%)\downarrow$ 1.31 2.02 2.22 39.65 66.16 60.34 18.29 15.10 23.61 25.85 31.35 83.04 84.75 88.24 56.67 54.85
-$M_{\text{non}}(\%)\downarrow$ 2.96 3.42 3.05 7.93 8.67 8.59 6.18 6.90 6.74 8.37 7.32 9.21 4.78 6.20 15.50 20.96
-Chinese-CLIP -L/14 $M_{\text{sem}}(\%)\uparrow$ 94.25 89.44 90.57 40.05 93.40 84.20 65.79 61.95 60.83 56.55 57.04 14.38 21.88 10.47 27.00 25.10
-$M_{\text{cul}}(\%)\downarrow$ 3.28 5.75 4.44 44.62 2.48 10.92 21.97 24.73 26.07 31.80 29.56 70.11 55.78 73.00 53.83 56.02
-$M_{\text{non}}(\%)\downarrow$ 2.46 4.81 4.99 15.32 4.13 4.88 12.24 13.32 13.10 11.65 13.40 15.51 22.34 16.54 19.17 18.89
-Jina-E-v4 $M_{\text{sem}}(\%)\uparrow$ 95.89 95.65 95.98 92.47 94.22 92.79 94.34 93.22 91.70 91.99 94.20 91.60 95.38 86.30 56.50 36.74
-$M_{\text{cul}}(\%)\downarrow$ 1.97 2.48 1.66 4.97 3.44 3.92 4.08 3.09 4.80 4.13 3.73 4.68 2.00 9.95 27.83 34.15
-$M_{\text{non}}(\%)\downarrow$ 2.13 1.86 2.36 2.55 2.34 3.29 1.58 3.69 3.50 3.88 2.07 3.72 2.62 3.75 15.67 29.11
-XLM-R-L/14 $M_{\text{sem}}(\%)\uparrow$ 89.66 91.15 89.46 90.32 92.30 89.40 87.63 85.37 86.12 87.86 87.98 90.95 93.07 89.41 80.33 40.49
-$M_{\text{cul}}(\%)\downarrow$ 4.11 4.04 5.83 4.44 4.13 6.26 6.58 6.54 7.39 7.16 5.80 4.85 3.54 3.75 8.50 24.71
-$M_{\text{non}}(\%)\downarrow$ 6.24 4.81 4.72 5.24 3.58 4.35 5.79 8.09 6.49 4.98 6.22 4.20 3.39 6.85 11.17 34.80
-XLM-R-B /16Plus $M_{\text{sem}}(\%)\uparrow$ 91.95 91.30 90.71 93.95 94.50 91.62 90.66 88.47 87.81 90.05 89.36 92.57 95.22 91.09 83.17 40.88
-$M_{\text{cul}}(\%)\downarrow$ 4.11 3.88 5.13 4.03 3.58 4.56 4.34 5.47 5.97 5.70 5.52 3.88 2.00 4.91 8.33 26.78
-$M_{\text{non}}(\%)\downarrow$ 3.94 4.81 4.16 2.02 1.93 3.82 5.00 6.06 6.23 4.25 5.11 3.55 2.77 4.01 8.50 32.34
-GME-Qwen2 -2B-Instruct $M_{\text{sem}}(\%)\uparrow$ 96.06 95.34 96.95 87.77 94.09 93.43 92.76 90.49 88.59 90.05 88.95 84.49 89.06 76.87 37.17 25.87
-$M_{\text{cul}}(\%)\downarrow$ 2.30 2.48 1.66 8.47 3.58 3.92 4.74 4.64 7.26 7.04 7.87 11.63 8.63 18.48 46.00 52.26
-$M_{\text{non}}(\%)\downarrow$ 1.64 2.17 1.39 3.76 2.34 2.65 2.50 4.88 4.15 2.91 3.18 3.88 2.31 4.65 16.83 21.86
-GME-Qwen2 -7B-Instruct $M_{\text{sem}}(\%)\uparrow$ 95.40 95.34 96.53 84.68 95.05 92.47 90.39 90.73 90.27 88.35 85.50 85.46 90.91 85.92 55.17 29.62
-$M_{\text{cul}}(\%)\downarrow$ 2.46 2.33 0.97 11.69 3.44 4.77 6.45 5.95 6.10 7.77 11.46 11.63 7.09 10.34 34.00 55.76
-$M_{\text{non}}(\%)\downarrow$ 2.13 2.33 2.50 3.63 1.51 2.76 3.16 3.33 3.63 3.88 3.04 2.91 2.00 3.75 10.83 14.62
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 SP $\downarrow$ 0.01 0.02 0.02 0.76 2.63 1.94 0.24 0.19 0.34 0.39 0.51 10.71 8.09 15.88 2.04 2.27
+TS $\downarrow$ 0.05 0.05 0.05 0.02 0.16 -0.05 -0.01 -0.02 -0.02 -0.01 -0.01 0.25 0.23 0.23 0.03 0.13
+IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.04 0.01 0.01 0.00 0.00 0.01 0.05 0.03 0.03 0.02 0.03
+CN-CLIP-L/14 SP $\downarrow$ 0.03 0.06 0.05 1.11 0.03 0.13 0.33 0.40 0.43 0.56 0.52 4.88 2.55 6.98 1.99 2.23
+TS $\downarrow$ -0.05 -0.05 -0.05 0.00 0.13 -0.07 -0.03 -0.03 -0.03 -0.01 -0.01 -0.40 0.93 -0.37 0.00 0.03
+IS $\downarrow$ 0.02 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.00 0.00 0.00 0.04 0.05 0.04 0.02 0.01
+Jina-E-v4 SP $\downarrow$ 0.02 0.03 0.02 0.05 0.04 0.04 0.04 0.03 0.05 0.04 0.04 0.05 0.02 0.12 0.49 0.93
+TS $\downarrow$ -0.01 -0.01 -0.01 -0.02 0.01 0.00 -0.01 -0.02 -0.02 -0.02 -0.02 0.02 0.01 0.00 0.00 0.13
+IS $\downarrow$ -0.01 0.00 -0.01 -0.01 0.00 0.01 0.00 0.00 -0.01 -0.01 -0.01 0.02 0.00 0.01 0.01 0.00
+XLM-R-L/14 SP $\downarrow$ 0.05 0.04 0.07 0.05 0.04 0.07 0.08 0.08 0.09 0.08 0.07 0.05 0.04 0.04 0.11 0.61
+TS $\downarrow$ -0.18 -0.18 -0.18 -0.11 -0.07 -0.10 -0.16 -0.16 -0.16 -0.12 -0.12 -0.07 -0.06 -0.17 -0.14 0.44
+IS $\downarrow$ 0.01 0.02 0.03 0.01 0.03 0.03 0.01 0.01 0.01 0.01 0.01 0.06 0.04 0.03 0.02 0.03
+XLM-R-B/16plus SP $\downarrow$ 0.05 0.04 0.06 0.04 0.04 0.05 0.05 0.06 0.07 0.06 0.06 0.04 0.02 0.05 0.10 0.66
+TS $\downarrow$ -0.24 -0.24 -0.24 -0.21 -0.19 -0.19 -0.24 -0.23 -0.23 -0.23 -0.23 -0.19 -0.19 -0.23 -0.23 0.49
+IS $\downarrow$ 0.01 0.01 0.00 0.00 0.01 0.01 0.00 0.00 -0.01 0.00 -0.01 0.03 0.02 0.03 0.02 0.00
+GME-Qwen2-2B SP $\downarrow$ 0.02 0.03 0.02 0.10 0.04 0.04 0.05 0.05 0.08 0.08 0.09 0.14 0.10 0.24 1.24 2.02
+TS $\downarrow$ 0.02 0.02 0.02 0.04 0.02 0.00 0.04 0.01 0.01 0.03 0.03 0.01 0.03 0.00 0.03 0.15
+IS $\downarrow$ 0.00 0.00 0.01 0.00 0.02 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.02 0.01 0.01
+GME-Qwen2-7B SP $\downarrow$ 0.03 0.02 0.01 0.14 0.04 0.05 0.07 0.07 0.07 0.09 0.13 0.14 0.08 0.12 0.62 1.88
+TS $\downarrow$ 0.04 0.04 0.04 0.05 0.02 0.02 0.07 0.02 0.02 0.04 0.04 0.09 0.11 0.03 0.04 0.14
+IS $\downarrow$ 0.00 0.00 0.01 0.00 0.01 0.01 0.00 0.00 0.00 0.00 0.00 0.03 0.02 0.01 0.01 0.01
+Table 12: Cross-Country and Cross-Model Comparison of Language and Cultural Bias Metrics.
+This table presents the results for the Self-Preference Cultural Bias score (SP), Text Silhouette (TS), and Image Silhouette (IS) scores across various multimodal retrievers for a selection of countries.
+Appendix O Detailed Results
+The full details of the RQ2 and RQ3 experiments, including all the win rates of all models, are illustrated in the Table 13 and 14, respectively.
+Model Metrics Country
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 $M_{\text{cr}}(\%)\uparrow$ 95.73 94.57 94.73 52.42 25.17 31.07 75.53 78.00 69.65 65.78 61.33 7.75 10.48 5.56 27.83 24.19
+$M_{\text{lb}}(\%)\downarrow$ 1.31 2.02 2.22 39.65 66.16 60.34 18.29 15.10 23.61 25.85 31.35 83.04 84.75 88.24 56.67 54.85
+$M_{\text{ti}}(\%)\downarrow$ 2.96 3.42 3.05 7.93 8.67 8.59 6.18 6.90 6.74 8.37 7.32 9.21 4.78 6.20 15.50 20.96
+CN-CLIP -L/14 $M_{\text{cr}}(\%)\uparrow$ 94.25 89.44 90.57 40.05 93.40 84.20 65.79 61.95 60.83 56.55 57.04 14.38 21.88 10.47 27.00 25.10
+$M_{\text{lb}}(\%)\downarrow$ 3.28 5.75 4.44 44.62 2.48 10.92 21.97 24.73 26.07 31.80 29.56 70.11 55.78 73.00 53.83 56.02
+$M_{\text{ti}}(\%)\downarrow$ 2.46 4.81 4.99 15.32 4.13 4.88 12.24 13.32 13.10 11.65 13.40 15.51 22.34 16.54 19.17 18.89
+XLM-R-L/14 $M_{\text{cr}}(\%)\uparrow$ 89.66 91.15 89.46 90.32 92.30 89.40 87.63 85.37 86.12 87.86 87.98 90.95 93.07 89.41 80.33 40.49
+$M_{\text{lb}}(\%)\downarrow$ 4.11 4.04 5.83 4.44 4.13 6.26 6.58 6.54 7.39 7.16 5.80 4.85 3.54 3.75 8.50 24.71
+$M_{\text{ti}}(\%)\downarrow$ 6.24 4.81 4.72 5.24 3.58 4.35 5.79 8.09 6.49 4.98 6.22 4.20 3.39 6.85 11.17 34.80
+XLM-R-B /16Plus $M_{\text{cr}}(\%)\uparrow$ 91.95 91.30 90.71 93.95 94.50 91.62 90.66 88.47 87.81 90.05 89.36 92.57 95.22 91.09 83.17 40.88
+$M_{\text{lb}}(\%)\downarrow$ 4.11 3.88 5.13 4.03 3.58 4.56 4.34 5.47 5.97 5.70 5.52 3.88 2.00 4.91 8.33 26.78
+$M_{\text{ti}}(\%)\downarrow$ 3.94 4.81 4.16 2.02 1.93 3.82 5.00 6.06 6.23 4.25 5.11 3.55 2.77 4.01 8.50 32.34
+ColQwen2.5 -3B-Multi $M_{\text{cr}}(\%)\uparrow$ 95.24 93.79 95.01 83.20 94.91 90.99 89.61 89.06 87.94 89.44 89.09 87.40 90.60 81.65 45.67 27.30
+$M_{\text{lb}}(\%)\downarrow$ 3.28 2.95 2.77 11.69 2.75 5.20 6.18 6.54 7.13 6.07 6.49 8.72 7.40 13.44 32.83 48.64
+$M_{\text{ti}}(\%)\downarrow$ 1.48 3.26 2.22 5.11 2.34 3.82 4.21 4.40 4.93 4.49 4.42 3.88 2.00 4.91 21.50 24.06
+ColQwen2.5 -7B-Multi $M_{\text{cr}}(\%)\uparrow$ 96.55 96.74 96.81 84.27 95.32 91.73 90.13 88.35 84.82 88.71 88.67 86.75 91.99 86.82 42.17 32.08
+$M_{\text{lb}}(\%)\downarrow$ 1.31 1.55 1.66 11.56 2.34 6.47 7.76 9.39 10.25 8.01 8.29 12.12 6.78 8.91 37.83 49.68
+$M_{\text{ti}}(\%)\downarrow$ 2.13 1.71 1.53 4.17 2.34 1.80 2.11 2.26 4.93 3.28 3.04 1.13 1.23 4.26 20.00 18.24
+ColQwen2.5 -v0.2 $M_{\text{cr}}(\%)\uparrow$ 93.10 89.60 92.09 83.60 94.09 88.55 90.79 87.51 86.64 90.29 90.61 85.14 88.60 79.59 38.83 29.88
+$M_{\text{lb}}(\%)\downarrow$ 4.11 4.97 3.47 9.95 3.16 6.68 5.00 7.37 6.87 5.22 4.83 8.72 8.94 13.82 42.83 42.95
+$M_{\text{ti}}(\%)\downarrow$ 2.79 5.43 4.44 6.45 2.75 4.77 4.21 5.11 6.49 4.49 4.56 6.14 2.47 6.59 18.33 27.17
+GME-Qwen2 -2B-Instruct $M_{\text{cr}}(\%)\uparrow$ 96.06 95.34 96.95 87.77 94.09 93.43 92.76 90.49 88.59 90.05 88.95 84.49 89.06 76.87 37.17 25.87
+$M_{\text{lb}}(\%)\downarrow$ 2.30 2.48 1.66 8.47 3.58 3.92 4.74 4.64 7.26 7.04 7.87 11.63 8.63 18.48 46.00 52.26
+$M_{\text{ti}}(\%)\downarrow$ 1.64 2.17 1.39 3.76 2.34 2.65 2.50 4.88 4.15 2.91 3.18 3.88 2.31 4.65 16.83 21.86
+GME-Qwen2 -7B-Instruct $M_{\text{cr}}(\%)\uparrow$ 95.40 95.34 96.53 84.68 95.05 92.47 90.39 90.73 90.27 88.35 85.50 85.46 90.91 85.92 55.17 29.62
+$M_{\text{lb}}(\%)\downarrow$ 2.46 2.33 0.97 11.69 3.44 4.77 6.45 5.95 6.10 7.77 11.46 11.63 7.09 10.34 34.00 55.76
+$M_{\text{ti}}(\%)\downarrow$ 2.13 2.33 2.50 3.63 1.51 2.76 3.16 3.33 3.63 3.88 3.04 2.91 2.00 3.75 10.83 14.62
+Jina-E-v4 $M_{\text{cr}}(\%)\uparrow$ 95.89 95.65 95.98 92.47 94.22 92.79 94.34 93.22 91.70 91.99 94.20 91.60 95.38 86.30 56.50 36.74
+$M_{\text{lb}}(\%)\downarrow$ 1.97 2.48 1.66 4.97 3.44 3.92 4.08 3.09 4.80 4.13 3.73 4.68 2.00 9.95 27.83 34.15
+$M_{\text{ti}}(\%)\downarrow$ 2.13 1.86 2.36 2.55 2.34 3.29 1.58 3.69 3.50 3.88 2.07 3.72 2.62 3.75 15.67 29.11
 Table 13: Cross-Country and Cross-Model Comparison of Win Percentages.
-This table presents the results for the Semantically Relevant, Culturally Relevant, and Non-Relevant Win Percentages across various multimodal retrievers for a selection of countries.
-Appendix K Computational Resource
-The experiment is performed with a single A100 GPU for approximately 3 gpu hours for each model or 54 hours in total with library version of colpali-engine 0.3.13.dev1+g9bee9b2b7, transformers 4.53.3 for most experiment except, GME models are inferenced under transformers 4.51.3
-Appendix L Authoring and Implementation Tools
+This table presents the results for the Correct, Language-biased, and Totally Irrelevant Win Percentages across various multimodal retrievers for a selection of countries.
+Model Metrics Country
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 $M_{\text{cr}}(\%)$ 80.39 82.17 79.94 43.36 4.12 35.23 50.33 48.28 40.86 37.14 36.80 0.81 0.76 1.03 13.79 17.16
+$M_{\text{orlb}}(\%)$ 8.33 6.05 7.47 33.56 51.92 31.12 28.52 32.94 42.28 38.83 40.36 51.05 58.38 53.81 42.86 23.35
+$M_{\text{or}}(\%)$ 7.68 8.06 8.71 3.62 3.71 4.64 8.28 11.30 8.04 8.86 7.52 1.13 0.76 0.26 2.49 5.03
+$M_{\text{cdr}}(\%)$ 3.10 3.72 3.60 10.07 3.16 12.76 7.88 4.64 5.06 6.80 6.84 0.81 0.00 0.90 8.97 20.65
+$M_{\text{lb}}(\%)$ 0.49 0.00 0.28 8.72 36.13 15.08 3.94 2.38 2.98 7.77 7.52 46.05 38.87 43.61 30.73 27.48
+$M_{\text{ti}}(\%)$ 0.00 0.00 0.00 0.67 0.96 1.16 1.05 0.48 0.78 0.61 0.96 0.16 1.22 0.39 1.16 6.32
+CN-CLIP -L/14 $M_{\text{cr}}(\%)$ 62.91 64.65 63.21 22.82 78.43 25.00 25.10 26.99 26.72 22.45 22.71 0.97 12.35 5.55 14.12 19.35
+$M_{\text{orlb}}(\%)$ 18.63 16.59 16.74 34.77 2.47 48.73 41.13 31.03 33.72 34.59 33.93 53.62 17.99 36.77 43.69 32.77
+$M_{\text{or}}(\%)$ 14.38 12.71 11.89 7.38 14.29 14.98 10.12 12.84 10.51 12.74 12.04 1.29 14.63 5.55 4.65 5.68
+$M_{\text{cdr}}(\%)$ 3.59 4.65 6.92 12.89 4.67 2.43 7.88 11.77 11.80 11.53 10.40 0.64 18.29 10.19 8.47 16.00
+$M_{\text{lb}}(\%)$ 0.33 0.62 0.55 17.85 0.00 7.81 11.70 12.60 13.10 14.08 14.36 42.83 19.82 34.06 25.58 20.00
+$M_{\text{ti}}(\%)$ 0.16 0.78 0.69 4.30 0.14 1.05 4.07 4.76 4.15 4.61 6.57 0.64 16.92 7.87 3.49 6.19
+XLM-R-L/14 $M_{\text{cr}}(\%)$ 80.56 82.64 79.81 78.93 75.69 73.84 78.98 77.41 80.16 75.97 76.74 74.24 76.52 69.16 69.93 32.90
+$M_{\text{orlb}}(\%)$ 8.17 6.05 6.50 9.53 9.48 8.76 8.28 10.34 9.73 11.89 13.27 7.25 8.54 9.03 4.65 6.19
+$M_{\text{or}}(\%)$ 6.70 6.36 8.99 8.32 9.34 9.39 7.62 7.73 5.06 8.74 5.88 10.95 9.76 14.19 12.46 9.94
+$M_{\text{cdr}}(\%)$ 4.58 4.96 4.01 2.55 4.67 7.49 4.60 4.40 4.93 3.40 3.56 6.44 4.88 5.29 9.47 33.29
+$M_{\text{lb}}(\%)$ 0.00 0.00 0.55 0.54 0.41 0.32 0.53 0.00 0.13 0.00 0.14 0.00 0.30 0.77 1.00 7.10
+$M_{\text{ti}}(\%)$ 0.00 0.00 0.14 0.13 0.41 0.21 0.00 0.12 0.00 0.00 0.41 1.13 0.00 1.55 2.49 10.58
+XLM-R-B /16Plus $M_{\text{cr}}(\%)$ 77.29 78.45 75.24 72.89 71.98 71.84 73.19 72.29 74.19 68.57 73.46 68.28 69.51 64.00 64.45 32.77
+$M_{\text{orlb}}(\%)$ 7.03 7.91 8.71 11.01 10.85 9.39 11.04 14.51 13.10 17.48 16.28 10.31 11.74 13.16 8.31 7.61
+$M_{\text{or}}(\%)$ 12.25 9.46 12.03 12.62 12.23 12.03 9.99 9.51 7.52 10.19 7.80 12.88 14.33 14.97 13.79 9.16
+$M_{\text{cdr}}(\%)$ 3.27 3.57 3.73 3.09 4.40 6.22 5.12 3.21 4.54 3.64 2.33 7.57 4.27 5.68 10.47 30.58
+$M_{\text{lb}}(\%)$ 0.16 0.31 0.28 0.40 0.27 0.11 0.26 0.24 0.39 0.12 0.14 0.48 0.15 1.16 1.50 8.65
+$M_{\text{ti}}(\%)$ 0.00 0.31 0.00 0.00 0.27 0.42 0.39 0.24 0.26 0.00 0.00 0.48 0.00 1.03 1.50 11.23
+ColQwen2.5 -3B-Multi $M_{\text{cr}}(\%)$ 70.42 72.25 70.95 54.23 69.51 57.81 54.27 56.48 54.22 54.25 55.95 62.96 49.85 39.74 26.08 23.35
+$M_{\text{orlb}}(\%)$ 13.24 11.32 13.00 28.32 15.11 27.95 32.19 28.42 29.57 27.79 29.69 17.71 36.59 35.74 34.72 28.39
+$M_{\text{or}}(\%)$ 15.03 13.02 14.11 14.77 11.40 11.71 10.78 13.20 13.10 15.66 11.35 14.01 9.45 15.10 6.98 3.35
+$M_{\text{cdr}}(\%)$ 0.98 2.95 1.52 1.61 3.30 1.90 2.10 1.31 1.69 1.46 2.19 4.03 1.68 4.39 14.95 21.29
+$M_{\text{lb}}(\%)$ 0.33 0.00 0.14 0.67 0.41 0.21 0.53 0.36 0.91 0.73 0.27 0.97 1.98 3.61 13.79 21.16
+$M_{\text{ti}}(\%)$ 0.00 0.47 0.28 0.40 0.27 0.42 0.13 0.24 0.52 0.12 0.55 0.32 0.46 1.42 3.49 2.45
+ColQwen2.5 -7B-Multi $M_{\text{cr}}(\%)$ 69.44 71.94 69.71 49.26 64.56 56.65 55.19 49.11 51.62 54.37 56.09 63.45 49.09 49.94 31.56 27.10
+$M_{\text{orlb}}(\%)$ 14.05 11.47 11.20 38.93 20.05 30.49 34.30 35.67 34.50 29.85 33.38 19.81 39.18 28.13 27.08 22.06
+$M_{\text{or}}(\%)$ 14.71 13.33 16.32 9.93 11.68 10.86 7.62 9.99 10.51 12.14 8.21 9.82 6.86 13.55 6.98 5.29
+$M_{\text{cdr}}(\%)$ 1.63 2.79 2.63 1.21 3.43 1.90 2.50 3.45 2.08 2.79 2.05 4.99 2.29 5.55 18.60 24.77
+$M_{\text{lb}}(\%)$ 0.00 0.00 0.00 0.54 0.14 0.11 0.39 1.31 1.30 0.36 0.00 1.77 2.29 1.68 13.12 17.81
+$M_{\text{ti}}(\%)$ 0.16 0.47 0.14 0.13 0.14 0.00 0.00 0.48 0.00 0.49 0.27 0.16 0.30 1.16 2.66 2.97
+ColQwen2.5 -v0.2 $M_{\text{cr}}(\%)$ 59.64 65.89 56.85 56.64 58.10 46.41 57.42 59.93 62.52 54.85 56.50 54.11 31.71 28.13 16.11 25.03
+$M_{\text{orlb}}(\%)$ 20.42 17.21 19.64 25.77 22.94 33.97 24.44 25.09 19.33 24.51 27.50 21.42 38.72 43.23 44.85 18.58
+$M_{\text{or}}(\%)$ 18.63 15.50 21.16 14.77 14.29 11.81 14.85 11.77 13.36 17.23 12.86 14.01 9.60 8.39 5.98 7.61
+$M_{\text{cdr}}(\%)$ 1.31 1.40 1.94 2.42 3.98 5.80 3.15 2.85 3.63 2.79 2.60 6.28 1.83 5.42 11.63 26.32
+$M_{\text{lb}}(\%)$ 0.00 0.00 0.00 0.27 0.41 1.79 0.13 0.24 0.52 0.24 0.41 3.54 17.68 13.42 19.27 15.87
+$M_{\text{ti}}(\%)$ 0.00 0.00 0.41 0.13 0.27 0.21 0.00 0.12 0.65 0.36 0.14 0.64 0.46 1.42 2.16 6.58
+GME-Qwen2 -2B-Instruct $M_{\text{cr}}(\%)$ 70.75 71.32 67.77 61.61 69.51 63.92 62.16 60.88 61.09 57.89 59.37 60.87 55.79 53.68 31.73 31.35
+$M_{\text{orlb}}(\%)$ 12.25 12.71 15.35 20.13 11.95 19.20 17.87 22.95 21.92 22.69 24.21 19.00 25.61 18.19 20.60 13.81
+$M_{\text{or}}(\%)$ 14.05 13.80 15.21 12.48 13.74 13.61 13.80 12.25 11.67 15.90 12.45 11.92 10.98 14.32 6.98 6.32
+$M_{\text{cdr}}(\%)$ 2.94 1.86 1.52 3.89 4.26 2.85 4.73 2.97 3.63 2.67 2.05 5.64 4.27 10.19 24.42 28.90
+$M_{\text{lb}}(\%)$ 0.00 0.00 0.00 0.94 0.41 0.21 1.05 0.48 1.17 0.49 1.50 1.29 2.13 2.45 11.79 13.29
+$M_{\text{ti}}(\%)$ 0.00 0.31 0.14 0.94 0.14 0.21 0.39 0.48 0.52 0.36 0.41 1.29 1.22 1.16 4.49 6.32
+GME-Qwen2 -7B-Instruct $M_{\text{cr}}(\%)$ 67.81 68.68 68.88 61.34 67.58 62.97 63.34 58.86 59.66 54.61 55.54 60.39 56.71 60.26 37.87 34.84
+$M_{\text{orlb}}(\%)$ 11.60 11.63 12.31 22.01 16.35 23.00 20.24 25.68 26.98 28.03 30.37 21.74 28.51 17.42 23.92 17.81
+$M_{\text{or}}(\%)$ 17.81 16.90 16.04 11.54 13.19 10.55 12.75 11.65 10.89 14.20 11.63 10.79 9.76 12.39 8.47 5.94
+$M_{\text{cdr}}(\%)$ 2.78 2.48 2.63 4.03 2.75 2.95 3.15 2.85 1.82 2.79 1.92 5.48 3.81 8.65 18.94 25.03
+$M_{\text{lb}}(\%)$ 0.00 0.16 0.00 0.67 0.14 0.42 0.39 0.71 0.52 0.36 0.55 1.13 0.91 0.90 7.81 11.48
+$M_{\text{ti}}(\%)$ 0.00 0.16 0.14 0.40 0.00 0.11 0.13 0.24 0.13 0.00 0.00 0.48 0.30 0.39 2.99 4.90
+Jina-E-v4 $M_{\text{cr}}(\%)$ 61.76 64.81 61.13 60.67 58.65 57.38 56.64 54.10 58.88 57.28 59.92 60.06 57.16 55.74 38.54 36.39
+$M_{\text{orlb}}(\%)$ 16.67 16.12 19.92 23.76 22.25 22.57 23.13 25.56 24.51 20.75 24.62 16.91 23.48 18.84 17.44 11.35
+$M_{\text{or}}(\%)$ 19.77 17.21 17.15 13.69 15.66 18.04 17.08 17.12 14.79 20.02 14.36 18.52 16.01 17.55 9.97 8.77
+$M_{\text{cdr}}(\%)$ 1.63 1.71 1.38 1.34 3.16 1.79 2.63 2.73 1.43 1.70 0.96 3.22 2.59 5.68 20.10 28.39
+$M_{\text{lb}}(\%)$ 0.00 0.00 0.00 0.27 0.14 0.11 0.13 0.24 0.39 0.12 0.14 0.64 0.46 1.55 8.64 9.55
+$M_{\text{ti}}(\%)$ 0.16 0.16 0.41 0.27 0.14 0.11 0.39 0.24 0.00 0.12 0.00 0.64 0.30 0.65 5.32 5.55
+Table 14: This table presents the results for the Correct, Object Relevant Language-biased, Object Relevant, Cultural Descriptor Relevant, and Totally Irrelevant Win Percentages across various multimodal retrievers for a selection of countries.
+Appendix P Analysis of Comparative Association Bias in Text-to-Image Retrieval
+To rigorously validate the failure modes and perspectival biases identified in RQ2 and RQ3, we employed a statistical evaluation targeting retrieval performance, embedding confusion, and the impact of query augmentation.
+P.1 Rank-1 Win Rate Analysis
+Complementing the continuous score analyses, we evaluated the model’s explicit retrieval preference at the top rank using a Chi-Square ($\chi^{2}$) Goodness of Fit Test.
+This approach determines whether the model exhibits a statistically significant tendency to select one specific failure mode over another when the ground truth is not prioritized.
+We defined the Null Hypothesis as the condition where the model is equally likely to retrieve either of the two competing distractor categories ($P(A)=P(B)$).
+To compute the statistic, the expected counts ($E$) for the two competing groups were derived by averaging their combined observed wins ($E_{A}=E_{B}=\frac{O_{A}+O_{B}}{2}$), while the count of unrelated or ground truth retrievals remained fixed as observed values.
+The resulting $\chi^{2}$ statistic, calculated with 1 degree of freedom, identifies whether the observed disparity in win rates is a result of systematic algorithmic bias.
+P.2 Test Pairs
+We conducted statistical analysis on specific pairs of candidate images, identified by the color codes in Figure 4, to isolate distinct levels of bias.
+The pairs were selected to evaluate how the model resolves conflicts between semantic content, explicit cultural descriptors, and implicit language priors.
+  • Query Language Association Bias Test (Object-Relevant Language Bias (Purple) vs. Object Relevant (Blue))
+    This comparison investigates association bias in scenarios where the model successfully retrieves the correct semantic object (e.g., a train) but fails to adhere to the explicit cultural descriptor.
+By comparing a candidate that aligns with the query’s language culture (Purple) against a random candidate (Blue), we test whether the script of the query exerts a stronger influence on visual selection than the explicit textual description.
+  • Cultural Descriptor Relevance (Yellow) vs. Language Bias (Pink):
+    This comparison analyzes the model’s prioritization mechanism during Semantic Failure.
+In cases where the model fails to retrieve the correct object entirely, we determine whether the retrieval is driven by a specific token match (e.g., the word "Japanese" leading to Japanese food; Yellow) or by the latent prior of the query language itself (e.g., Thai script leading to Thai cultural imagery; Pink).
+This reveals whether the error stems from specific keyword fixation or broad language-modality leakage.
+P.3 Statistical Test Results
+Table 15 summarizes the overall results of the statistical significance tests.
+For a breakdown of each test, Tables 16 and 17 provide the detailed $p$-values for each model across all evaluated countries.
+In the Query Language Association Bias Test, the majority of architectures, specifically the Vision-Language Contrastive models and LLM-Based Retrieval Embedding models, exhibited a statistically significant preference for the Object-Relevant Language Bias candidate (Purple) over the neutral alternative (Blue).
+This consistent rejection of the null hypothesis for these models confirms the presence of a strong association bias, demonstrating that the script of the input query affects the retrieved image result.
+However, the Cross-lingual Alignment models (XLM-R) served as an exception, showing no statistically significant difference ($p>0.05$) between the candidates, suggesting these models are less susceptible to implicit script bias when object relevance remains constant.
+Conversely, the Cultural Descriptor vs.
+Language Bias comparison reveals a divergence in model behavior during semantic failure.
+The LLM-Based Retrieval Embeddings models and Cross-lingual Alignment models displayed a statistically significant preference for the Cultural Descriptor Relevance candidate (Yellow), indicating a general tendency to prioritize specific explicit tokens (e.g., "Japanese") over the implicit language prior.
+Notably, standard CLIP-L/14 exhibited a statistically significant preference for the Language-biased category (Pink) over the Cultural Descriptor (Diff <0<0), suggesting that within CLIP’s embedding space, the implicit bias derived from the query language script outweighs the explicit semantic cultural tokens.
+Model QL Association CD vs QL
+(Blue vs Purple) (Yellow vs Pink)
+Blue Purple Diff $p$-val Sig. Yellow Pink Diff $p$-val Sig.
+Vision-Language Contrastive Models
+CLIP-L/14 0.053 0.298 -0.245 0.000 True 0.085 0.126 -0.040 0.000 True
+CN-CLIP-L/14 0.098 0.245 -0.147 0.000 True 0.100 0.112 -0.012 0.006 True
+Cross-lingual Alignment Models
+XLM-R-L/14 0.068 0.072 -0.004 0.290 False 0.076 0.003 0.073 0.000 True
+XLM-R-B/16plus 0.097 0.099 -0.002 0.647 False 0.066 0.004 0.062 0.000 True
+LLM-Based Retrieval Embedders
+ColQwen2.5-3B-M 0.093 0.207 -0.114 0.000 True 0.094 0.013 0.081 0.000 True
+ColQwen2.5-7B-M 0.078 0.177 -0.099 0.000 True 0.098 0.011 0.087 0.000 True
+ColQwen2.5-v0.2 0.126 0.201 -0.076 0.000 True 0.096 0.021 0.074 0.000 True
+GME-Qwen2-2B 0.115 0.184 -0.069 0.000 True 0.081 0.016 0.065 0.000 True
+GME-Qwen2-7B 0.108 0.208 -0.100 0.000 True 0.068 0.014 0.054 0.000 True
+Jina-E-v4 0.136 0.196 -0.060 0.000 True 0.071 0.012 0.059 0.000 True
+Table 15: Win rates, differences, and statistical significance for the Object Relevance and Cultural Descriptor studies.
+Results are significant if $p<0.05$.
+(Blue is the Object Relevant category, matching only the object.
+Purple represents the Object Relevant Language-biased category, which matches both the object and the culture of the language of the text query.
+Yellow represents the Cultural Descriptor Relevant category, which matches the CD’s culture, and Pink is the Language-biased category, which matches only the culture of the language of the text query.)
+Model Metrics Country
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 Diff -0.01 0.01 0.02 -0.18 -0.45 -0.23 -0.19 -0.14 -0.23 -0.28 -0.36 -0.49 -0.58 -0.53 -0.21 -0.09
+pp-val 0.60 0.43 0.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+CN-CLIP-L/14 Diff 0.00 -0.02 -0.01 -0.14 0.06 -0.08 -0.23 -0.14 -0.13 -0.17 -0.22 -0.41 -0.19 -0.39 -0.25 -0.07
+pp-val 1.00 0.30 0.74 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Jina-E-v4 Diff 0.00 -0.01 -0.08 -0.12 -0.01 -0.05 -0.09 -0.05 -0.10 -0.05 -0.10 -0.02 -0.07 -0.05 -0.06 -0.07
+pp-val 0.94 0.73 0.00 0.00 0.65 0.01 0.00 0.03 0.00 0.02 0.00 0.39 0.00 0.01 0.00 0.00
+XLM-R-L/14 Diff 0.02 0.01 0.03 0.01 0.00 0.01 0.00 0.00 -0.04 -0.06 -0.08 0.01 0.00 0.02 0.02 0.01
+pp-val 0.20 0.33 0.03 0.32 0.85 0.70 0.77 0.72 0.01 0.00 0.00 0.38 0.77 0.21 0.17 0.39
+XLM-R-B/16plus Diff 0.02 0.04 0.02 0.02 0.03 0.02 -0.01 -0.04 -0.07 -0.08 -0.04 0.04 0.00 0.03 0.01 0.01
+pp-val 0.24 0.02 0.26 0.27 0.06 0.28 0.42 0.02 0.00 0.00 0.03 0.02 0.80 0.14 0.48 0.17
+ColQwen2.5-3B-M Diff -0.03 0.01 0.00 -0.16 -0.03 -0.16 -0.22 -0.15 -0.15 -0.13 -0.20 -0.06 -0.18 -0.18 -0.10 -0.01
+pp-val 0.07 0.69 0.79 0.00 0.14 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.51
+ColQwen2.5-7B-M Diff 0.02 0.00 0.03 -0.13 -0.06 -0.11 -0.23 -0.17 -0.15 -0.11 -0.16 -0.05 -0.17 -0.12 -0.09 -0.01
+pp-val 0.21 0.92 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.45
+ColQwen2.5-v0.2 Diff -0.05 -0.03 -0.02 -0.05 -0.02 -0.10 -0.13 -0.05 -0.06 -0.03 -0.13 -0.01 -0.16 -0.15 -0.18 -0.04
+pp-val 0.05 0.13 0.28 0.01 0.33 0.00 0.00 0.01 0.00 0.12 0.00 0.76 0.00 0.00 0.00 0.01
+GME-Qwen2-2B Diff 0.02 0.01 0.00 -0.07 0.00 -0.09 -0.09 -0.11 -0.09 -0.07 -0.18 -0.10 -0.08 -0.06 -0.13 -0.03
+pp-val 0.41 0.64 0.89 0.00 0.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.03
+GME-Qwen2-7B Diff 0.05 0.03 0.02 -0.12 -0.01 -0.11 -0.14 -0.14 -0.11 -0.12 -0.24 -0.15 -0.14 -0.09 -0.16 -0.13
+pp-val 0.03 0.13 0.29 0.00 0.50 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Table 16: Win rate differences and $p$-values for each model across different countries in the QL Association Test.
+Model Metrics Country
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 Diff 0.04 0.04 0.04 0.06 -0.24 0.07 0.05 0.04 0.04 0.03 0.01 -0.46 -0.38 -0.41 0.10 0.23
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.67 0.00 0.00 0.00 0.00 0.00
+CN-CLIP-L/14 Diff 0.05 0.05 0.05 0.08 0.05 0.01 0.04 0.06 0.04 0.03 0.02 -0.39 -0.24 -0.34 0.05 0.19
+pp-val 0.00 0.00 0.00 0.00 0.00 0.22 0.00 0.00 0.00 0.06 0.26 0.00 0.00 0.00 0.05 0.00
+Jina-E-v4 Diff 0.02 0.02 0.02 0.04 0.05 0.05 0.04 0.03 0.02 0.03 0.02 0.04 0.02 0.12 0.18 0.25
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.00
+XLM-R-L/14 Diff 0.05 0.05 0.06 0.04 0.04 0.05 0.04 0.04 0.05 0.05 0.04 0.06 0.06 0.05 0.09 0.40
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+XLM-R-B/16plus Diff 0.03 0.04 0.05 0.04 0.04 0.05 0.04 0.03 0.03 0.03 0.03 0.08 0.03 0.05 0.08 0.35
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+ColQwen2.5-3B-M Diff 0.08 0.10 0.10 0.07 0.03 0.02 0.06 0.03 0.04 0.04 0.04 0.09 0.02 0.07 0.20 0.34
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.12 0.00 0.00 0.00
+ColQwen2.5-7B-M Diff 0.04 0.07 0.07 0.08 0.04 0.02 0.07 0.05 0.05 0.07 0.07 0.07 0.02 0.11 0.23 0.35
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00
+ColQwen2.5-v0.2 Diff 0.05 0.04 0.05 0.12 0.06 0.03 0.08 0.05 0.06 0.07 0.05 0.11 0.02 0.10 0.07 0.22
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.00 0.00 0.00
+GME-Qwen2-2B Diff 0.02 0.04 0.03 0.06 0.04 0.03 0.05 0.03 0.03 0.06 0.01 0.07 0.04 0.11 0.14 0.28
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.00 0.00 0.00 0.00 0.00
+GME-Qwen2-7B Diff 0.03 0.04 0.04 0.05 0.03 0.02 0.05 0.03 0.04 0.06 0.02 0.05 0.02 0.07 0.13 0.20
+pp-val 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.02 0.00 0.00 0.00 0.00 0.00
+Table 17: Win rate differences and $p$-values for each model across different countries in the CD vs QL test.
+Appendix Q Similarity Drift
+Q.1 Similarity Drift Significance
+Table 18 presents the detailed quantitative breakdown of the Similarity Drift ($\Delta\text{sim}$) analysis across all evaluated languages.
+In general, adding a cultural descriptor results in a positive similarity shift across most models, validating that explicit cultural context helps narrow the retrieval scope toward the target culture.
+The data reveals a critical failure mode in the baseline CLIP-L/14 model regarding non-Latin scripts.
+While CLIP performs robustly for Western languages (e.g., English, German), it exhibits negligible drift for Chinese and Thai.
+This negative value implies that for certain script-heavy or linguistically complex contexts in standard CLIP, adding a cultural descriptor introduces embedding noise rather than semantic clarity, effectively pushing the retrieved result further away from the ground truth.
+In contrast, multilingual models such as XLM-R and GME-Qwen2 maintain consistent positive drift across all tested regions, confirming their superior cross-cultural alignment capabilities.
+Model Difference by Query Language Culture
+US GB AU DE CN JP FR ES AR PT BR SA TH IN KE NG
+CLIP-L/14 4.89 4.54 4.68 4.93 0.64 3.43 3.55 4.34 4.51 3.14 3.02 0.21 0.18 -0.16 2.81 3.74
+CN-CLIP-L/14 5.04 5.09 5.36 4.40 5.27 2.80 3.58 4.44 4.52 3.52 3.59 -0.51 0.01 0.02 3.20 5.61
+GME-Qwen2-2B 7.47 6.41 7.51 7.21 7.78 7.08 7.30 8.31 9.36 7.32 7.84 7.26 8.03 6.96 6.14 8.15
+GME-Qwen2-7B 3.91 3.62 3.98 3.48 7.81 6.29 1.60 3.41 4.01 2.10 2.03 3.77 4.96 2.17 1.69 5.36
+Jina-E-v4 2.82 2.52 2.57 1.71 0.73 2.05 1.67 2.15 2.66 1.72 1.28 2.29 2.69 1.90 2.76 1.62
+XLM-R-L/14 5.71 5.28 5.31 4.13 3.70 3.06 4.96 4.92 5.15 4.93 4.87 3.86 3.97 2.82 4.64 5.65
+XLM-R-B/16Plus 9.52 9.10 9.34 6.57 5.55 4.40 8.82 8.05 8.20 8.18 7.98 5.41 5.91 3.92 8.42 11.77
+Table 18: Mean Similarity Drift ($\Delta Sim\times 100$) by Model and Query Language Culture.
+The table measures the change in cosine similarity to the Correct when a specific cultural descriptor is added to the query.
+Positive values indicate successful instruction following, while near-zero or negative values (bolded) indicate that the model fails to process the cultural descriptor or treats it as noise.
+Appendix R Association Bias Evaluation in Verbose Text Query.
+We extended the association bias evaluation to encompass scenarios both in the absence of and incorporating explicit cultural descriptors (in both RQ2 and RQ3) in the case where we increase the verbosity of the prompt without giving more object or cultural cues.
+R.1 Association Bias in Verbose Queries without Cultural Descriptors
+Our framework evaluates association bias in detailed text queries using two primary formats: task-based and casual.
+The former specifies the intended use of the image (e.g., “a picture of a train for my homework”), whereas the latter captures the user’s situational intent (e.g., “I am looking for a bicycle”).
+Query generation was facilitated by Gemini [^3] Version used: gemini-2.5-flash (Released September 25, 2025).
+using the prompt templates illustrated in Figure 21.
+The evaluation result is shown in Figure 19.
+The results indicate that the accuracy for verbose text queries across both task-based and casual formats decreases by 6–9% compared to short queries (RQ2) for CLIP-L/14 and CN-CLIP-L/14.
+In contrast, Cross-lingual Alignment Models and MLLM-based retrieval models remain stable, with deviations of less than 4%, except for ColQwen2.5-7B-M, which decreases by 6% in the task-based format.
+The drop in accuracy occurs because the additional words in the text can distract the model from the core concept of the query.
+This issue is more severe in CLIP-L/14 and CN-CLIP-L/14, as these models appear to struggle with non-Latin script languages (with the exception of Chinese and Japanese for CN-CLIP-L/14).
+Consequently, as the query length increases, the model becomes more biased toward the culture associated with the language rather than the object.
+R.2 Association Bias in Verbose Queries with a Cultural Descriptor
+To evaluate association bias in verbose queries with a cultural descriptor, a country descriptor is added to the neutral text queries from subsection R.1 using Gemini2, following the prompt shown in Figure 20.
+An example of a resulting query is “A picture of a Japanese train for my homework”.
+The evaluation result is shown in Figure 20, short queries consistently achieve the highest correct answer across almost all models.
+The additional words in query cause models to drift away from the target concept or culture, especially in CLIP and CN-CLIP models, where accuracy drops by approximately 9-12% in verbose formats.
+This performance loss is primarily driven by an increase in the “Object Relevant, Language-biased” category, indicating that verbose queries cause these models to over-rely on the language’s cultural priors.
+Furthermore, a distinct “Language Bias” failure mode (Pink) is observed in CLIP models, which worsens with verbosity, confirming a tugging effect toward pure language heuristics.
+In contrast, the “Cultural Diversity Relevant” category (Yellow) remains stable (5-10%) for most architectures but diminishes in CLIP models, suggesting that the additional text dilutes the model’s focus on explicit cultural descriptors.
+Figure 19: Comparative association bias across verbose text query in task-based and casual format without cultural descriptors.
+Figure 20: Comparative association bias across verbose text query in task-based and casual format with cultural descriptors.
+Figure 21: The prompt given to Gemini to generate unique country-specific image concepts.
+**Role:** You are a Senior Computational Linguist and Polyglot with native-level fluency in the primary languages of the
+following countries:
+China (ZH), India (HI), Japan (JA), Saudi Arabia (AR), France (FR), Germany (DE), Brazil (PT-BR), Kenya (SW), Thailand (TH),
+USA (EN-US), Spain (ES-ES), Argentina (ES-AR), UK (EN-GB), Australia (EN-AU), Nigeria (EN-NG), and Portugal (PT-PT).
+**Task:**
+1. **Modify:** Take the English sentences provided in ‘format_task‘ and ‘format_casual‘. Insert a country-specific
+reference (e.g., concept: train, mention country: China, original sentence: I need images of a Train for a school project.
+The modified sentence: I need images of a chinese train for a school project.) into the English sentence.
+2. **Translate:** Translate the modified sentences into the native language of the "Target Country."
+**Strict Constraints:**
+1. **Grammar & Fluency:** Adjust the sentence structure, classifiers (e.g., Thai ’khan’), and gender markers to ensure
+the sentence is natural for a native speaker in that country.
+2. **No Stereotypes:** Keep the sentences "Culture Agnostic." Do not add descriptors like "high-speed" for China or
+"vintage" for India. Only add the country name/adjective.
+3. **Target Language Mapping:**
+    *   Nigeria: Yoruba
+**Input Data:**
+‘‘‘json
+  "concept": "bus",
+  "format_task": "I need an image of a Bus for my school project.",
+  "format_casual": "I’m just looking for some cool Bus pictures to check out."
+**Output Format:**
+Return ONLY a valid JSON object following this exact structure:
+‘‘‘json
+  "bus": {
+    "format_task": {
+      "Nigeria": {
+        "mention_China": "Translated string here...",
+        "mention_India": "Translated string here...",
+        "mention_Japan": "...",
+        "mention_Saudi Arabia": "...",
+        "mention_France": "...",
+        "mention_Germany": "...",
+        "mention_Brazil": "...",
+        "mention_Kenya": "...",
+        "mention_Thailand": "...",
+        "mention_USA": "...",
+        "mention_Spain": "...",
+        "mention_Argentina": "...",
+        "mention_UK": "...",
+        "mention_Australia": "...",
+        "mention_Nigeria": "...",
+        "mention_Portugal": "..."
+    "format_casual": {
+      "Nigeria": {
+        "mention_China": "Translated casual string here...",
+Figure 22: The prompt given to Gemini to generate unique country-specific image concepts.
+Appendix S Semantically Focused Instruction Prompt Effects
+This section investigates the impact of system prompts on MLLM retriever performance.
+We evaluate three distinct prompts: (1) a baseline default prompt, (2) a standard text-to-image retriever prompt ("Find an image that matches the given caption"), and (3) a culturally agnostic prompt ("Find an image that matches the given caption.
+Focus on the semantics instead of the textual language").
+The results are presented in Figure 23.
+As shown, the associative bias remains relatively consistent across different prompt configurations.
+This suggests that varying the system prompt does not significantly alter the model’s retrieval behavior.
+Figure 23: Comparative association bias across default prompt, standard text-to-image retriever prompt (T2I), and a culturally agnostic prompt (Cul-Ag).
+Appendix T Computational Resource
+The experiment is performed with a single A100 GPU for approximately 3 gpu hours for each model or 54 hours in total with library version of colpali-engine 0.3.13.dev1+g9bee9b2b7, transformers 4.53.3 for most experiments except GME models are utilized under transformers 4.51.3
+Appendix U Authoring and Implementation Tools
 In preparing this manuscript, we utilized several generative large language models.
-For language editing and stylistic refinement, we employed Google’s Gemini 2.5-flash, along with models from xAI’s Grok family (e.g., Grok-3 Expert and Fast variants).
-For assistance with code implementation, scripting, and debugging, we used a model from Anthropic’s Claude series (e.g., Claude 4.0 Sonnet).
-Appendix M Detailed Results
-The full details of RQ2 experiment including all win rate of all models are illustrated in the Table 11.
-Appendix N 3XCM Dataset Benckmark Samples
-This research provides an association evaluation benchmark and image metadata. Examples of the benchmark and image metadata are shown in Figure 18 and Figure 19, respectively.
-Figure 18: Examples of 3XCM dataset benchmark
-Figure 19: Metadata for image of 3XCM dataset benchmark
+For language editing and stylistic refinement, we employed Google’s Gemini, along with models from xAI’s Grok family (e.g., Grok Expert and Fast variants).
+For assistance with code implementation, scripting, and debugging, we used a model from Anthropic’s Claude series (e.g., Claude Sonnet).